### v1

# Value Residual Learning For Alleviating Attention Concentration In Transformers

Abstract
Transformers can capture long-range dependencies using self-attention, allowing tokens to attend to all others directly. However, stacking multiple attention layers leads to attention concentration. One natural way to address this issue is to use cross-layer attention, allowing information from earlier layers to be directly accessible to later layers. However, this approach is computationally expensive. To address this problem, we propose Transformer with residual value (ResFormer) which approximates cross-layer attention through adding a residual connection from the values of the the first layer to all subsequent layers. Based on this method, one variant is the Transformer with single layer value (SVFormer), where all layers share the same value embedding from first layer, reducing the $KV$ cache by nearly 50%. Comprehensive empirical evidence demonstrates that ResFormer mitigates attention concentration problem in deeper layers and enhances representation across most layers, outperforming the vanilla Transformer, DenseFormer, and NeuTRENO in training error as well as downstream tasks. SVFormer trains significantly faster than the vanilla Transformer and performs better than other methods like GQA and CLA, with performance influenced by sequence length and cumulative learning rate.  [^1] Code is available at https://github.com/Zcchill/Value-Residual-Learning.
1 Introduction
The Transformer (Vaswani, 2017) model has become one of the leading architectures in recent years, excelling in both language modeling (Devlin, 2018, Lan, 2019, Brown, 2020) and computer vision tasks (Dosovitskiy, 2020). The discovery of scaling laws (Hoffmann et al., 2022, Kaplan et al., 2020) has driven the pursuit of larger Transformer models by increasing network depth and width.
Training large models presents significant challenges. Balancing the depth and width of a Transformer model within a fixed parameter budget is particularly difficult. While research indicates that deeper models generalize more compositionally than shallower ones (Petty et al., 2024), the training and deployment of deep models remain problematic. Although Transformers use residual connections (He et al., 2016) to address the vanishing gradient issue, training very deep Transformers is still challenging. For example, a 32-layer Vision Transformer (ViT) may perform worse than a 24-layer one (Zhou et al., 2021). This is mainly due to the smoothing mechanism of attention (Shi et al., 2022), which can lead to an over-smoothing effect (Nguyen et al., 2023) where the token representations become the same as the model’s depth increases.
Existing solutions to alleviate the over-smoothing problem in Transformer include adding extra regularizers (Nguyen et al., 2023, Shi et al., 2022) and optimizing the information flow within the model (Pagliardini et al., 2024). During the era of convolutional neural network architectures, Stochastic Depth (Huang et al., 2016) reduces the likelihood of over-smoothing by randomly dropping layers during training and DenseNet (Huang et al., 2017) mitigates the impact of over-smoothing by allowing each layer to directly access the hidden states of all preceding layers. Recently, DenseFormer (Pagliardini et al., 2024) adopts the idea of DenseNet when training Transformer. Additionally, NeuTRENO (Nguyen et al., 2023) alleviates over-smoothing through incorporating the difference between the value vectors of the first layer and the current layer to the attention output.
In this paper, we address the problem of multi-layer attention from a novel perspective. We introduce the phenomenon of attention concentration, which describes how a model’s attention increasingly focuses on fewer tokens as the network depth increases, as illustrated in Fig. 1 (Right). We quantify the degree of attention concentration using the entropy of the distribution of token importance, where lower entropy indicates a more pronounced concentration. In a deep model, the attention mechanism exhibits a pattern of “concentration - dispersion - concentration,” where the model could potentially lose useful information during the concentrated phase. See Fig. 15 for analysis of over-smoothing.
To alleviate the attention concentration from stacking multiple attention layers, an effective method is to use cross-layer attention on information from earlier layers. Given the high computational cost of cross-layer attention, we propose ResFormer as an approximation. ResFormer achieves a similar effect by applying a residual connection between the value vectors of the current layer and the first layer before the attention operation. Experimental results show that ResFormer mitigates the attention concentration effect in deeper layers, maintaining attention dispersion throughout the sequence, as shown in Fig. 1 (Middle).
During inference, deep networks require substantial $KV$ cache, severely impacting model deployment (Xiao et al., 2023). Existing $KV$-efficient methods often process keys and values simultaneously. Building on ResFormer, we decouple the value from the attention operation and propose a new kind of Transformer with single layer value (SVFormer). In SVFormer, the queries and keys of all layers share the value from the first layer. Consequently, SVFormer can save nearly half of the $KV$ cache. Additionally, SVFormer is orthogonal to the classical method GQA, and they can be used concurrently. Experimental results show that SVFormer trains significantly faster than the vanilla Transformer. Its performance mainly depends on factors like training sequence length and cumulative learning rate.
We experiment on a 20B SlimPajama sub-sampled dataset, using settings similar to popular large language models (Wei et al., 2023, Dubey et al., 2024, Kaplan et al., 2020). We compare different models by their relative training curves against the vanilla Transformer. Results show that ResFormer outperforms the vanilla Transformer, DenseFormer, and NeuTRENO across all settings. Furthermore, SVFormer trains faster than the vanilla Transformer and performs better when the sequence length is longer.
Figure 1: (Left) Illustration of the relative training loss (loss of target model - loss of vanilla Transformer) curve between different Transformer variants; model size is fixed to be 82M. (Middle) The average entropy of token importance across layers in ResFormer vs. the vanilla Transformer, where token importance is derived from the attention matrix. Lower entropy indicates more focused attention on specific tokens. More details can be found in Eqn. 11. (Right) The average entropy of token importance across layers in Llama (8B) (Dubey et al., 2024) and Mistral (7B) (Jiang et al., 2023).
(a) Transformer
(b) NeuTRENO
(c) DenseFormer
(d) ResFormer
(e) SVFormer
Figure 2: Simplified illustration of the vanilla Transformer, NeuTRENO, DenseFormer, ResFormer, and SVFormer, with only three-layer structures and no operations other than attention. $A^{i}$, $V^{i}$, and $H^{i}$ denote the attention matrix, value vectors, and attention outputs at the $i$-th layer, respectively. $\oplus$, $\ominus$, and $\otimes$ represent standard matrix addition, subtraction, and multiplication, respectively.
2 Related Work
2.1 Shortcut Connections For Better Information Flow
Deep learning models often consist of multiple layers, posing a challenge to minimize information loss during transmission. ResNet (He et al., 2016) mitigates the vanishing gradient problem with identity connections. Stochastic Depth (Huang et al., 2016) enhances training by randomly dropping layers. DenseNet (Huang et al., 2017) allows subsequent layers to directly access the hidden states of all preceding layers. These two methods further enhance the information flow after ResNet.
Related research indicates that for advanced Transformer architectures, although increasing depth continues to yield performance improvements in language modeling tasks, the gains become less significant with further increases (Petty et al., 2024). Furthermore, Zhou et al. (2021) illustrates that a 32-layer ViT underperforms a 24-layer ViT. Depth-Wise Attention (ElNokrashy et al., 2024) allows each query to access the key and value at the same position from previous layers through an attention-like mechanism before the output layer. DenseFormer (Pagliardini et al., 2024) integrates weighted fusion of outputs from all preceding layers after each layer. To explore why increasing depth in Transformers does not yield expected gains, Wang et al. (2022) finds that self-attention acts as a low-pass filter, smoothing token representations in ViTs. Additionally, Shi et al. (2022) investigates over-smoothing from a graph perspective in BERT-based language modeling tasks. NeuTRENO (Nguyen et al., 2023) adds the difference between the value vectors of the first and current layers to each layer’s attention output and significantly alleviates the over-smoothing problem.
In contrast to these methods, ResFormer accesses and integrates information from previous layers prior to the attention operation, as illustrated in Fig. 2. Moreover, it does not require the selection or tuning of additional hyperparameters.
2.2 $KV$ cache compressing
The $KV$ cache is a key factor limiting the efficiency of long-text model inference. Research in this area can be broadly classified into Transformer-based methods, which target redundant information in Transformer models, and non-Transformer methods, which mainly addresses the quadratic time complexity of attention with respect to sequence length.
For non-Transformer methods, Mamba (Gu & Dao, 2023) and RWKV (Peng et al., 2023) are two popular works. They replace the original softmax-based attention with SSM (Gu et al., 2021) and AFT (Zhai et al., 2021) mechanisms, respectively. Besides, several approaches have been proposed to enhance models’ ability to process long texts while reducing the reliance on $KV$ cache. Dai (2019) advocates segmenting long texts into smaller parts for attention computation. Building upon this, Munkhdalai et al. (2024) proposes using a fixed-size memory matrix for storing and retrieving historical information.
Transformer-based methods can be categorized into three main groups. The first group consists of post-training methods like SnapKV (Li et al., 2024) and ThinK (Xu et al., 2024), which compress KV cache during inference based on attention matrices at token or hidden dimension levels. The second group focuses on quantization and adopts low-precision $KV$ cache quantization rather than completely eliminating them (Hooper et al., 2024). The third group aims to maximize the efficiency of attention-based models via parameter or activation value sharing. The most representative works include Multi-Query Attention (Shazeer, 2019) and Grouped-Query Attention (Ainslie et al., 2023) which suggest to share key and value across a group of queries. MLKV (Zuhri et al., 2024) further suggest to share keys and values for queries across layers and MLA (Liu et al., 2024) introduces low-rank projection when processing keys and values. Besides, CLA (Brandon et al., 2024) and LISA (Mu et al., 2024) respectively point out that we can reuse keys, values, or the attention matrix across layers to reduce redundancy between layers. While these methods typically process both key and value simultaneously, SVFormer is the first approach to decouple value from query and key. It shares values across all layers and is compatible with other methods like GQA.
3 Method
3.1 Motivation: Information Transfer Via Cross Layer Attention
Let $\mathbf{H}_{n}\in\mathbb{R}^{l\times d}$ be the input hidden state of the $n$-th layer, where $l$ denotes the sequence length and $d$ is the dimension size. In standard attention, the hidden state $\mathbf{H}_{n}$ will be firstly projected into $\mathbf{Q}_{n},\mathbf{K}_{n},\mathbf{V}_{n}\in\mathbb{R}^{l\times d}$ through three linear projections $\mathbf{W^{\mathbf{Q}}},\mathbf{W^{\mathbf{K}}},\mathbf{W^{\mathbf{V}}}\in\mathbb{R}^{d\times d}$ respectively. For simplicity, we introduce dot-product attention of layer $n$ as
$\operatorname{Attention}(\mathbf{Q}_{n},\mathbf{K}_{n},\mathbf{V}_{n})$ $=\operatorname{Softmax}(\frac{\mathbf{Q}_{n}\mathbf{K}_{n}^{T}}{\sqrt{d}})\mathbf{V}_{n}.$ (1)
An ideal way to incorporate previous layers’ information is cross layer attention. The attention mechanism naturally extracts relevant information from previous layers. If these layers contain low-quality information, the similarity between the current layer’s query and the previous layers’ keys will be low, thus minimizing negative impacts. Given $m<n$ and the information $(\mathbf{Q}_{m},\mathbf{K}_{m},\mathbf{V}_{m})$ of $m$-th layer, the cross layer mechanism calculates the attention output $\mathbf{U}_{n}$ of $n$-th layer by the following attention formula:
$\mathbf{U}_{n}$ $=\operatorname{Softmax}\Big{(}{\mathbf{Q}_{n}\operatorname{Concat}(\mathbf{K}_{n},\mathbf{K}_{m})^{T}}/{\sqrt{d}}\Big{)}\operatorname{Concat}(\mathbf{V}_{n},\mathbf{V}_{m}).$ (2)
In practice, cross-layer attention enhances feature fusion by allowing information to flow between layers, capturing both intra-layer and inter-layer dependencies. However, this approach introduces additional computational overhead due to the concatenation of keys and values from multiple layers. For example, in scenarios described by Eqn. 2, the overall computational complexity of the model nearly doubles compared with vanilla attention described in Eqn. 1.
3.2 Efficient Cross Layer Attention
To solve this problem, we propose to replace the $\mathbf{K}_{m}$ with $\mathbf{K}_{n}$ in Eqn. 2, as shown in Eqn. 3.
$\mathbf{U}_{n}$ $\approx\operatorname{Softmax}\Big{(}{\mathbf{Q}_{n}\operatorname{Concat}(\mathbf{K}_{n},\mathbf{K}_{n})^{T}}/{\sqrt{d}}\Big{)}\operatorname{Concat}(\mathbf{V}_{n},\mathbf{V}_{m})$ (3)
$=\frac{1}{2}\operatorname{Softmax}\Big{(}{\mathbf{Q}_{n}\mathbf{K}_{n}^{T}}/{\sqrt{d}}\Big{)}({\mathbf{V}_{n}}+{\mathbf{V}_{m}}).$ (4)
Utilizing the concept of block matrices, Eqn. 3 can be further simplified into Eqn. 4. This simplification converts the concatenation operation of the two value matrices into an addition operation. Compared to Eqn. 1, this new method only brings a minimal increase in computational complexity while still leveraging the information from the $m$-th layer in the $n$-th layer. Furthermore, Eqn. 4 can be generalized to incorporate cross-layer attention across all preceding $n-1$ layers as follows:
$\mathbf{U}_{n}$ $\approx\frac{1}{n}\mathbf{A}_{n}\sum_{i=1}^{n}{\mathbf{V}_{i}}.$ (5)
Figure 3: Average token similarity between the outputs of different mapping methods and that of Eqn. 2.
Figure 4: Ablation study on sharing keys or values in every two layers, with CLAttention denoting sharing both.
where $\mathbf{A}_{n}$ denotes the original attention matrix for layer $n$. From the perspective of information propagation, model described by Eqn. 3 projects the historical values into the current layer’s embedding space using the current layer’s attention as a weight matrix. For example, a naive approach would be to perform identity mapping, as described by
$\vspace{-5mm}\mathbf{U}_{n}$ $=\mathbf{A}_{n}\mathbf{V}_{n}+\frac{1}{n-1}\sum_{i=1}^{n-1}{\mathbf{V}_{i}}.$ (6)
To evaluate the approximation effect of replacing the $\mathbf{K}_{m}$ with $\mathbf{K}_{n}$, we randomly select 1,000 pre-training data samples. For each layer of a trained baseline model, assuming cross-layer attention defined by Eqn. 2 is required for each layer with respect to the previous one, we calculate the cosine similarity between the outputs from Eqn. 2 and Eqn. 3. We also calculate the cosine similarity between the outputs from Eqn. 2 and Eqn. 6 for comparison. Fig. 4 shows that using current attention as a mapping matrix provides a better approximation for cross-layer attention.
3.3 Transformer with Residual Value
Based on Eqn. 5, we propose a variant of Transformer with residual value (ResFormer) which only chooses first layer as the target of cross layer attention since the first layer contains all basic information of each token. The analysis of entropy in Fig. 1 (Right) supports this point, indicating that attention tends to be relatively dispersed across different tokens in the initial layers of the model. The attention mechanism of ResFormer can be formulated as
$\mathbf{U}_{n}$ $=\frac{1}{2}\mathbf{A}_{n}({\mathbf{V}_{n}}+{\mathbf{V}_{1}}).$ (7)
where $n\geq 2$ and standard attention is applied in the first layer. From the training perspective, it explicitly learns a residual mapping instead of directly learning the desired underlying mapping and that’s why we call it ResFormer.
3.4 A unified View Of NeuTRENO and DenseFormer
Using our framework, the NeuTRENO can be defined as
$\mathbf{U}_{n}$ $=\Big{(}\mathbf{A}_{n}-\bm{\lambda}\mathbf{J}\Big{)}{\mathbf{V}_{n}}+\bm{\lambda}{\mathbf{V}_{1}}.$ (8)
where $\mathbf{J}$ denotes a matrix of ones and $\bm{\lambda}$ is a hyper-parameter. It can be found that the term of $\bm{\lambda}\mathbf{J}$ may have certain negative impact on the learning of original attention. If we ignore the attention output projections and the MLP layers, DenseFormer can also be modeled within our framework as
$\mathbf{U}_{n}$ $=\sum_{i=1}^{n}{\bm{\alpha}_{i}\mathbf{A}_{i}\mathbf{V}_{i}}.$ (9)
where $\{\bm{\alpha}_{i}\}_{i=1}^{n}$ is a set of hyper-parameters. DenseFormer uses attention matrix of previous layer as the weight matrix of projecting values but this is not aligned with the concept of the efficient cross layer attention shown in Eqn. 3.
3.5 SVFormer: Single-layer Value For Half $KV$ Cache
After ResFormer, a natural idea is whether we can remove the value vectors in each layer and have all layers share the value vectors from the first layer. We call this method SVFormer. Similar to ResFormer, SVFormer still adopts standard attention in the first layer and obtain the attention output $\mathbf{U_{n}}$ for $n$-th layer where $n\geq 2$ through
$\mathbf{U}_{n}$ $=\mathbf{A}_{n}{\mathbf{V}_{1}}.$ (10)
Compared to previous methods, SVFormer is the first method that decouple value vectors from attention. Its main advantage is that it only requires computing and storing the value vectors for the first layer, saving nearly half of the $KV$ cache during inference. Similar methods like CLA reduce $KV$ cache by sharing both of the key and value vectors every two layers, but the results in Fig. 4 show that sharing values has less negative impact on performance compared with sharing keys.
4 Pretrain Experiments
4.1 Setting
4.1.1 Training Details
Following Brandon et al. (2024), we choose the Llama-like architecture and SlimPajama (Soboleva et al., 2023) data for main experiments. Specifically, the architecture includes pre-normalization, SwiGLU activations (Shazeer, 2020), rotary position embedding (Su et al., 2024), and no dropout. For slimpajama, we randomly sample nearly 20B tokens according to the original data distribution of seven domains during training and adopt tokenizer used for “RedPajama-INCITE-7B-Base”. The details of training data can be found in Table 2 in Appendix.
Unless otherwise noted, we train all models using AdamW optimizer with 0.1 weight decay, $\beta_{1}=0.9$, $\beta_{2}=0.95$ and the max grad norm 1.0. The batch size is set to be around 2M tokens (Zhang et al., 2024) with a sequence length of 2,048 and the total steps is fixed 10,000 steps (Kaplan et al., 2020). We adopt linear learning rate warmup for the first 1,200 steps with the initial learning rate and the peak learning rate to be 1e-7 and 6e-4 respectively. The cosine decay schedule gradually decays to 10% of the peak learning rate by the end of training (Brandon et al., 2024, Zhou et al., 2024, Wei et al., 2023). See Table 4 and Table 3 in Appendix for more details.
All models are trained with 8 Nvidia A100 80G GPUs using mixed-precision training in FP16. We adopt deepspeed zero-2 optimizer and flash attention mechanism.
Figure 5: (Left) The relative training curve between a 82M ResFormer and Transformer across different training sequence lengths. (Middle) Average training loss for the final 50 steps across different model sizes and the corresponding fitted curves. (Right) The relative training curve across different model size for a fixed 2,048 training sequence length.
4.1.2 Relative Training Loss Curve On SlimPajama
We trained all models for only one epoch on SlimPajama subsets, and primarily use training loss to compare different models. Furthermore, we use the relative training loss curve for better visualizing the difference among different models from the perspective of loss landscape. Specifically, for each method, we will subtract the smoothed training curve of the vanilla Transformer, obtained under the same experimental settings, from the smoothed training curves of the method. The smoothing is done using a window size of 10 steps or 100 steps.
4.1.3 Entropy For Analyzing Attention Concentration Effects
Given the attention matrix $\mathbf{A}\in\mathbb{R}^{l\times l}$ at one layer, we use entropy $\bm{e}$ to represent its concentration effect. To obtain entropy $\bm{e}$, calculate the importance vector $\mathbf{a}=\frac{1}{l}\sum_{j=1}^{l}A_{ij}$ firstly where $\mathbf{A}$ is a lower triangular matrix. The entropy can be formulated as follows:
$\bm{e}$ $=-\sum_{i=1}^{l}\bm{a}_{i}^{\prime}\log{\bm{a}^{\prime}_{i}}.$ (11)
where $a_{i}^{{}^{\prime}}={a_{i}}/\Big{(}{\sum_{i=1}^{l}{a_{i}}}\Big{)}$ for $i=1,2,\dots,l$ and the higher the entropy $\bm{e}$, the greater the degree of clustering in $\bm{a}$, i.e., attention matrix $\mathbf{A}$ is more likely to focus on several specific tokens. 1,000 examples with a sequence length of 2,048 in the training dataset are selected for entropy analysis.
4.1.4 Spectral Decomposition For Analyzing Representations
Spectral Decomposition is a classical method to analyze the representations of models. Zhu et al. (2021) suggests that the eigenvectors with larger eigenvalues are more transferable. Here we use spectral decomposition to analyze the feature space of value $\bm{v}$ of one layer as following:
$\frac{1}{l}\sum_{i=1}^{l}\bm{v}_{i}\bm{v}_{i}^{T}$ $=\sum_{i=j}^{d}\bm{w}_{j}\bm{\lambda}_{j}\bm{w}_{j}^{T}.$ (12)
where $\bm{w}_{j}$ is the $j$-th eigenvector with eigenvalue $\bm{\lambda}_{j}$ for $i=1,2,\dots,d$ and $d$ is the dimensionality of the value’s feature space.
4.2 ResFormer vs. vanilla Transformer
We trained ResFormer and vanilla Transformer with different model size on data with different sequence lengths. In Fig. 5, ResFormer consistently outperforms vanilla Attention throughout training for different models sizes. Additionally, the results in Fig. 1 (Left) illustrate that ResFormer outperforms DenseFormer and NeuTRENO. Furthermore, integrating ResFormer with NeuTRENO leads to additional performance improvements.
Figure 6: Ablation study of adding residual connection to queries or keys.
Figure 7: Ablation study of adding residual connection using different mapping matrix.
Figure 8: Ablation studies on which historical layer’s value to include in residual connections.
4.3 Ablation study of residual connection
In Eqn. 4, we employ residual connections for the values. We compare this approach with models that add residual connections to queries or keys. The results, shown in Fig. 8, indicate that only residual connections for values yield positive effects. One possible explanation is that attention mechanisms are sensitive to perturbations, and modifying queries or keys can significantly impact attention matrix.
Moreover, we compare with the models based on Eqn. 2 and Eqn. 6. From the cross layer attention perspective, the former uses a gold attention matrix in Eqn. 2 to map historical values, while the latter uses an all-ones matrix in Eqn. 6. The results in Fig. 8 align with Fig. 4, showing that identity mapping causes significant perturbations, leading to poor performance. Interestingly, using current attention as the mapping matrix results in an even lower final loss than using gold attention.
When determining the mapping method and target value, it is crucial to consider which historical layers’ values should be included in the residual connection. Fig. 8 shows that each Transformer layer should add a shortcut to the first layer’s value rather than to the nearest preceding layer or all previous layers, highlighting the first-layer value’s critical importance. A potential explanation is that incorporating values from other layers may dilute the impact of the first-layer value.
4.4 Downstream Evaluations
We compare the different models on several classical reasoning tasks following (Zhang et al., 2024) in a zero-shot way. The tasks include Hellaswag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2019), ARC-Easy and ARC-Challenge (Clark et al., 2018) and PIQA (Bisk et al., 2020). The results in Table 1 show that ResFormer (82M) achieves an average accuracy improvement of nearly 3% compared to the vanilla Transformer (82M).
Model Max Length HellaSwag Obqa WinoGrande ARC-c ARC-e PIQA Avg.
Transformer 2,048 0.263 0.142 0.492 0.199 0.331 0.572 0.333
ResFormer 2,048 0.273 0.148 0.512 0.182 0.414 0.604 0.355
Transformer 64,000 0.267 0.142 0.485 0.179 0.322 0.570 0.328
ResFormer 64,000 0.274 0.136 0.513 0.184 0.407 0.588 0.350
Table 1: Zero-shot accuracy on commonsense reasoning tasks.
4.5 Visualization of ResFormer
Figure 9: Left: Distribution of eigenvalues for the value vectors in the first layer of ResFormer and Transformer. Right: Maximum eigenvalue for each layer of ResFormer and Transformer.
To figure out why ResFormer can achieve better performance on language modeling tasks than vanilla Transformer, we conduct visualization based on the eigenvalue decomposition discussed in Section 4.1.4. After sorting the eigenvalues in descending order, we compute the average eigenvalue for each layer across 1,000 randomly sampled pre-train data examples. The results in Fig. 9 indicate that the value vectors generated by most layers of the ResFormer exhibit stronger representational capacity compared to those of the vanilla Transformer. Besides, we analyze the attention concentration effects mentioned in Section 4.1.3 using the same batch of test data. Fig. 1 (Middle) illustrates that the clustering effect of attention increases significantly with the number of layers for the vanilla Transformer, whereas the clustering effect is relatively less pronounced for the ResFormer.
4.6 SVFormer vs. GQA
Figure 10: The relative training loss for SVFormer and other $KV$ efficient model compared with vanilla attention. The numbers in parentheses represent the training sequence length. Left: Model with nearly ${1}/{2}$ $KV$ cache. Right: Model with nearly ${1}/{8}$ $KV$ cache.
In the Fig. 10, at a training sequence length of 64,000, SVFormer demonstrates lower final loss compared to existing KV-efficient methods such as CLA and GQA. Moreover, it can be used concurrently with GQA to enhance KV efficiency further. However, we observed that with a training sequence length of 2,048, SVFormer underperforms compared to GQA. The results indicate that sequence length significantly affects SVFormer’s performance. Thus, we conducted more comprehensive experiments on sequence length.
Results in Fig. 11 (Left) demonstrate that SVFormer will always be gradually surpassed by vanilla attention during training while its training speed is much faster than vanilla attention. However, as the training sequence length increases, the SVFormer model performs better. In this way, we focus on the critical point, defined as the number of training steps exceeded. Fig. 11 (Right) illustrates that the relationship between the critical point and sequence length exhibits an exponential trend. We argue that it’s due to the challenge deep models face in fully optimizing the increasingly larger first-layer value matrix as the training sequence length grows.
Figure 11: Left: The relative training loss for SVFormer under different sequence lengths with a fixed batch size of 2M tokens. Right: Analysis of critical point, and we predict it for length 64,000 using linear regression with the last 1,000 data points.
4.7 Other Factors Influencing SVFormer
Intuitively, the training effectiveness of SVFormer is influenced by factors such as the maximum learning rate, warmup steps, model size, and other factors beyond just the training sequence length. We conducted experiments to explore these relationships.
Based on the results shown in Fig. 12(a) and Fig. 12(b), a smaller learning rate benefits SVFormer more, with warmup’s impact being comparatively small. This could be attributed to the model’s outcomes being closely tied to the total summed learning rate, which has weak connection with warmup steps (Kaplan et al., 2020). Moreover, larger models often require smaller learning rates to ensure training stability, making them more suitable for using SVFormer.
Except for the 2M model, Llama-like models ranging from 82M to 468M, as well as models with the GPT2 architecture, exhibit similar critical points and final losses (see Fig. 12(c) and Fig. 12(d)). This suggests that the difference between SVFormer and the vanilla Transformer is not sensitive to model size and architecture.
(a) Learning Rate.
(b) Warmup Steps.
(c) Model Size.
(d) Architecture.
Figure 12: The relative training loss for SVFormer under different hyper-parameter setting.
4.8 Ablation study of SVFormer
Figure 13: Ablation study of sharing first layer’s query(key) across all layers.
Figure 14: Ablation study on sharing values from different numbers of layers.
To better understand SVFormer, we conduct several ablation experiments. We first observe the effects of sharing the first layer’s queries or keys across all layers in Fig. 14, finding that this significantly impacts model performance, similar to the results in Fig. 4. Additionally, sharing the first layer’s values in a multi-layer network may reduce the network’s “effective depth.” By updating the shared values using intermediate layers as “anchors,” we find that increasing the number of “anchors” improves performance, as shown in Fig. 14.
5 Conclusion
In this paper, we propose the concept of attention concentration, a problem that arises from stacking multiple attention layers. From the perspective of cross-layer attention, we derive ResFormer, which adds a residual connection between the value vectors of the current layer and those of the first layer before the attention operation to alleviate attention concentration. Additionally, we introduce SVFormer, based on ResFormer, which reduces the $KV$ cache by nearly half. We conducted comprehensive experiments on the language modeling task to validate the advantages of these two Transformer variants in different scenarios.
Acknowledgement
This work was supported by the Research Center for Industries of the Future at Westlake University (Grant No. WU2023C017) and the Key Research.
Ethics Statement
On the one hand, the data employed in this paper is sourced from publicly available datasets provided by the company, which have undergone a certain level of filtering. On the other hand, the models trained in our study are solely utilized for experimental analysis and will not be publicly deployed.
Reproducibility Statement
We have detailed the complete experiments setup such as batch size, optimizer and learning rates in Section 4.1.1. Besides, we have released the source codes. These resources should be sufficient to reproduce the results of the paper.
Appendix A Appendix
A.1 Token Similarity Analysis
Attention concentration tends to make embeddings of different tokens more similar, resulting in over-smoothing. The extent of over-smoothing can be assessed by calculating the average token similarity $\bm{s}$ of the hidden states using the following formula:
$\bm{s}$ $=\frac{1}{l(l-1)}\sum\nolimits_{i=1}^{l}\sum\nolimits_{j=1,j\neq i}^{l}\operatorname{Sim}\Big{(}\bm{h}_{i},\bm{h}_{j}\Big{)}.$ (13)
where $\{\bm{h}_{i}\}_{i=1}^{l}$ is the hidden state of the $i$-th token and $\operatorname{Sim}(\cdot)$ denotes the operation of cosine similarity. The results in Fig. 15 are align with the results in Fig. 1. In the case of Llama and Mistral, the average token similarity demonstrates an “M”-shaped pattern with increasing network depth, while entropy follows a “W”-shaped pattern at corresponding positions. These trends suggest that attention concentration indeed leads to the phenomenon of over-smoothing.
Figure 15: (Left) The average token similarity of hidden states across layers in ResFormer vs. the vanilla Transformer. (Right) The average token similarity of hidden states across layers in Llama (8B) (Dubey et al., 2024) and Mistral (7B) (Jiang et al., 2023).
A.2 Pre-train Dataset
Data source proportions Tokens
Commoncrawl 50% 10 B
C4 20% 4 B
GitHub 10% 2 B
Books 5% 1 B
ArXiv 5% 1 B
Wikpedia 5% 1 B
StackExchange 5% 1 B
Table 2: The details of pre-train dataset.
Based on the equation $D\geq 5000\cdot N^{0.74}$ (Kaplan et al., 2020) where $D$ is data size and $N$ is the number of non-embedding parameters, we need to collect at least 17.5B for model has N = 700M non-embedding parameters (corresponding to complete 1B model with 2,048 hidden size, 50,277 vocab size and 2,048 sequence length) to avoid over-fitting. Besides, Xie et al. (2024) indicates that the mixture proportions of pre-training data domains significantly affects the training results. In this way, we sampled 20B tokens data from original 627B data based on the original data proportions shown in the Table 2.
A.3 Training Details
Max Sequence Length 512 2,048 8,192 32,000 64,000
Total Batch Size 4,096 1,024 256 64 32
Per-GPU Batch Size 128 32 8 2 1
Gradient Accumulation Step 32
GPUs 8
Table 3: Training details for training dataset with different sequence length.
Section 4.1.1 introduces the main experimental hyperparameters used in the paper. This section further details the training parameters for various model sizes and training sequence lengths. Table 4 demonstrates the differences among models of various sizes. The configurations for the number of layers, attention heads, hidden dimensions, and FFN dimensions are based on Biderman et al. (2023). Additionally, the $\bm{\lambda}$ in Eqn. 8 is set to be 0.4 for NeuTRENO. Moreover, as reported in Table 3, the batch size that a single GPU can accommodate varies depending on the length of the training sequences. Note that the total number of tokens in each batch is consistently 2 million.
Model Size 2M 82M 180M 468M
Layers 4 8 12 24
Attention Heads 2 8 12 16
Hidden Dimension 16 512 768 1,024
FFN Dimension 56 1,792 2,688 3,584
Tie Word Embedding False
(Peak Learning Rate, Final Learning Rate) $(6e-4,6e-5)$
Learning Rate Schedule Cosine Decay
Vocabulary Size 50,277
Activation Function SwiGLU
Position Embedding RoPE ($\theta$ = 10,000)
Batch Size 2M tokens
Data Size 20B tokens
(Warmup Steps, Training Steps) (120, 10,000)
Adam $\beta$ (0.9, 0.95)
Dropout 0.0
Weight Decay 0.1
Table 4: Training details for models with different size.
A.4 Validation Loss On Slimpajama
Section 4.1.2 introduces to use relative training loss as a main evaluation matrix. Table 5 reports the validation loss for differnt model on the whole validation split of slimpajama.
Model
Common
Crawl
C4 Github
Stack
Exchange
Wikipedia Book Arxiv Avg.
Transformer (82M) 3.3595 3.5388 1.4247 2.3872 2.9047 3.3797 2.1779 2.7389
Transformer (180M) 3.0961 3.2834 1.2451 2.1651 2.5897 3.1309 2.0001 2.5015
Transformer (468M) 2.8514 3.0430 1.0908 1.9628 2.2821 2.8979 1.8362 2.2806
Resformer (82M) 3.3362 3.5191 1.3941 2.3592 2.8646 3.3572 2.1518 2.7117
Resformer (180M) 3.0631 3.2504 1.2200 2.1350 2.5435 3.0994 1.9732 2.4692
Resformer (468M) 2.8214 3.0115 1.0730 1.9388 2.2477 2.8696 1.8142 2.2537
Table 5: Validation loss on slimpajama.

### v2

@@
 Comprehensive empirical evidence demonstrates that ResFormer mitigates attention concentration problem in deeper layers and enhances representation across most layers, outperforming the vanilla Transformer, DenseFormer, and NeuTRENO in training error as well as downstream tasks.
+Further visualization results suggest that Resformer alleviates attention sinks through avoiding value-state drains.
 SVFormer trains significantly faster than the vanilla Transformer and performs better than other methods like GQA and CLA, with performance influenced by sequence length and cumulative learning rate.
@@
 Recently, DenseFormer (Pagliardini et al., 2024) adopts the idea of DenseNet when training Transformer.
-Additionally, NeuTRENO (Nguyen et al., 2023) alleviates over-smoothing through incorporating the difference between the value vectors of the first layer and the current layer to the attention output.
+Additionally, NeuTRENO (Nguyen et al., 2023) alleviates over-smoothing through incorporating the difference between the value states of the first layer and the current layer to the attention output.
 In this paper, we address the problem of multi-layer attention from a novel perspective.
@@
 Given the high computational cost of cross-layer attention, we propose ResFormer as an approximation.
-ResFormer achieves a similar effect by applying a residual connection between the value vectors of the current layer and the first layer before the attention operation.
+ResFormer achieves a similar effect by applying a residual connection between the value states of the current layer and the first layer before the attention operation.
 Experimental results show that ResFormer mitigates the attention concentration effect in deeper layers, maintaining attention dispersion throughout the sequence, as shown in Fig.
 1 (Middle).
+Further visualization results show that Resformer effectively alleviates attention sinks (Xiao et al., 2023) through avoiding value-state drains (Guo et al., 2024b).
 During inference, deep networks require substantial $KV$ cache, severely impacting model deployment (Xiao et al., 2023).
@@
 Figure 2: Simplified illustration of the vanilla Transformer, NeuTRENO, DenseFormer, ResFormer, and SVFormer, with only three-layer structures and no operations other than attention.
-$A^{i}$, $V^{i}$, and $H^{i}$ denote the attention matrix, value vectors, and attention outputs at the $i$-th layer, respectively.
+$A^{i}$, $V^{i}$, and $H^{i}$ denote the attention matrix, value states, and attention outputs at the $i$-th layer, respectively.
 $\oplus$, $\ominus$, and $\otimes$ represent standard matrix addition, subtraction, and multiplication, respectively.
@@
 (2022) investigates over-smoothing from a graph perspective in BERT-based language modeling tasks.
-NeuTRENO (Nguyen et al., 2023) adds the difference between the value vectors of the first and current layers to each layer’s attention output and significantly alleviates the over-smoothing problem.
+NeuTRENO (Nguyen et al., 2023) adds the difference between the value states of the first and current layers to each layer’s attention output and significantly alleviates the over-smoothing problem.
 In contrast to these methods, ResFormer accesses and integrates information from previous layers prior to the attention operation, as illustrated in Fig.
@@
 For example, a naive approach would be to perform identity mapping, as described by
-$\vspace{-5mm}\mathbf{U}_{n}$ $=\mathbf{A}_{n}\mathbf{V}_{n}+\frac{1}{n-1}\sum_{i=1}^{n-1}{\mathbf{V}_{i}}.$ (6)
+$\mathbf{U}_{n}$ $=\mathbf{A}_{n}\mathbf{V}_{n}+\frac{1}{n-1}\sum_{i=1}^{n-1}{\mathbf{V}_{i}}.$ (6)
 To evaluate the approximation effect of replacing the $\mathbf{K}_{m}$ with $\mathbf{K}_{n}$, we randomly select 1,000 pre-training data samples.
@@
 3.5 SVFormer: Single-layer Value For Half $KV$ Cache
-After ResFormer, a natural idea is whether we can remove the value vectors in each layer and have all layers share the value vectors from the first layer.
+After ResFormer, a natural idea is whether we can remove the value states in each layer and have all layers share the value states from the first layer.
 We call this method SVFormer.
@@
 $\mathbf{U}_{n}$ $=\mathbf{A}_{n}{\mathbf{V}_{1}}.$ (10)
-Compared to previous methods, SVFormer is the first method that decouple value vectors from attention.
-Its main advantage is that it only requires computing and storing the value vectors for the first layer, saving nearly half of the $KV$ cache during inference.
-Similar methods like CLA reduce $KV$ cache by sharing both of the key and value vectors every two layers, but the results in Fig.
+Compared to previous methods, SVFormer is the first method that decouple value states from attention.
+Its main advantage is that it only requires computing and storing the value states for the first layer, saving nearly half of the $KV$ cache during inference.
+Similar methods like CLA reduce $KV$ cache by sharing both of the key and value states every two layers, but the results in Fig.
 4 show that sharing values has less negative impact on performance compared with sharing keys.
@@
 $\bm{e}$ $=-\sum_{i=1}^{l}\bm{a}_{i}^{\prime}\log{\bm{a}^{\prime}_{i}}.$ (11)
-where $a_{i}^{{}^{\prime}}={a_{i}}/\Big{(}{\sum_{i=1}^{l}{a_{i}}}\Big{)}$ for $i=1,2,\dots,l$ and the higher the entropy $\bm{e}$, the greater the degree of clustering in $\bm{a}$, i.e., attention matrix $\mathbf{A}$ is more likely to focus on several specific tokens.
-1,000 examples with a sequence length of 2,048 in the training dataset are selected for entropy analysis.
+where $\bm{a}_{i}^{\prime}={\bm{a}_{i}}/\Big{(}{\sum_{i=1}^{l}{\bm{a}_{i}}}\Big{)}$ for $i=1,2,\dots,l$ and the higher the entropy $\bm{e}$, the greater the degree of clustering in $\bm{a}$, i.e., attention matrix $\mathbf{A}$ is more likely to focus on several specific tokens.
 4.1.4 Spectral Decomposition For Analyzing Representations
@@
 In Fig.
-5, ResFormer consistently outperforms vanilla Attention throughout training for different models sizes.
+5, ResFormer consistently outperforms vanilla Transformer throughout training for different models sizes.
 Additionally, the results in Fig.
@@
 Furthermore, integrating ResFormer with NeuTRENO leads to additional performance improvements.
+We further test the variant of Resformer defined as $\mathbf{U}_{n}=\mathbf{A}_{n}({\mathbf{V}_{n}}+{\bm{\lambda}}{\mathbf{V}_{1}})$.
+As shown in Fig.6, Resformer can accommodate a wide range of $\bm{\lambda}$ values and the performance improves as $\bm{\lambda}$ increases, achieving the best results at ${\bm{\lambda}}=2$.
+Regardless of the value of $\bm{\lambda}$, ResFormer consistently outperforms Transformers.
+It suggests that the success of Resformer lies in the use of $\mathbf{V}_{1}$ and the transformation through $\mathbf{A}_{n}$.
+The ablation study of different hyperparameters $\bm{\lambda}$ for NeuTRENO, as defined in Equation 8, can be found in the Appendix A.3.
+[Uncaptioned image] Figure 6: Ablation study of different $\bm{\lambda}$ for Resformer.
 Figure 7: Ablation study of adding residual connection to queries or keys.
@@
 4.3 Ablation study of residual connection
+(a) Token importance.
+(b) Norms of value states.
+(c) Norms of hidden states.
+Figure 10: The token importance, value-state norms, and hidden-state norms of the first token in each sequence across layers in ResFormer vs. the vanilla Transformer.
+(a) Transformer.
+(b) NeuTRENO.
+(c) Resformer.
+Figure 11: The distribution of token importance for different models at different layers.
 In Eqn.
@@
 The results, shown in Fig.
-8, indicate that only residual connections for values yield positive effects.
+9, indicate that only residual connections for values yield positive effects.
 One possible explanation is that attention mechanisms are sensitive to perturbations, and modifying queries or keys can significantly impact attention matrix.
@@
 Table 1: Zero-shot accuracy on commonsense reasoning tasks.
+Figure 12: Left: Distribution of eigenvalues for the value states in the first layer of ResFormer and Transformer. Right: Maximum eigenvalue for each layer of ResFormer and Transformer.
 4.5 Visualization of ResFormer
-Figure 9: Left: Distribution of eigenvalues for the value vectors in the first layer of ResFormer and Transformer. Right: Maximum eigenvalue for each layer of ResFormer and Transformer.
 To figure out why ResFormer can achieve better performance on language modeling tasks than vanilla Transformer, we conduct visualization based on the eigenvalue decomposition discussed in Section 4.1.4.
@@
 The results in Fig.
-9 indicate that the value vectors generated by most layers of the ResFormer exhibit stronger representational capacity compared to those of the vanilla Transformer.
+12 indicate that the value states generated by most layers of the ResFormer exhibit stronger representational capacity compared to those of the vanilla Transformer.
 Besides, we analyze the attention concentration effects mentioned in Section 4.1.3 using the same batch of test data.
@@
 1 (Middle) illustrates that the clustering effect of attention increases significantly with the number of layers for the vanilla Transformer, whereas the clustering effect is relatively less pronounced for the ResFormer.
+We further visualize the attention weights, value-state norms $\|\bm{v}\|_{2}$, and hidden-state norms $\|\bm{h}\|_{2}$ of tokens at different layers and positions, with detailed results in Appendix A.2.
+Given that attention clustering often occurs on the first token, we primarily show its results in Fig.
+10.
+The results indicate that using ResFormer significantly mitigates attention sinks (Xiao et al., 2023), value-state drains (Guo et al., 2024b) and residual-state peaks (Sun et al., 2024).
+Guo et al.
+(2024a) attributes these phenomena to the mutual reinforcement mechanism of model and we suggest that the value shortcut disrupts this mechanism by alleviating value-state drains.
+Specifically, for tokens lacking semantic information, such as the first token, a large value state magnitude can adversely affect the prediction of future tokens if it is overly attended to, resulting in higher loss.
 4.6 SVFormer vs. GQA
-Figure 10: The relative training loss for SVFormer and other $KV$ efficient model compared with vanilla attention.
-The numbers in parentheses represent the training sequence length.
-Left: Model with nearly ${1}/{2}$ $KV$ cache.
-Right: Model with nearly ${1}/{8}$ $KV$ cache.
 In the Fig.
-10, at a training sequence length of 64,000, SVFormer demonstrates lower final loss compared to existing KV-efficient methods such as CLA and GQA.
+13, at a training sequence length of 64,000, SVFormer demonstrates lower final loss compared to existing KV-efficient methods such as CLA and GQA.
 Moreover, it can be used concurrently with GQA to enhance KV efficiency further.
@@
 We argue that it’s due to the challenge deep models face in fully optimizing the increasingly larger first-layer value matrix as the training sequence length grows.
+Figure 13: The relative training loss for SVFormer and other $KV$ efficient model compared with vanilla attention.
+The numbers in parentheses represent the training sequence length.
+Left: Model with nearly ${1}/{2}$ $KV$ cache.
+Right: Model with nearly ${1}/{8}$ $KV$ cache.
 Figure 14: Left: The relative training loss for SVFormer under different sequence lengths with a fixed batch size of 2M tokens.
@@
 Based on the results shown in Fig.
-12(a) and Fig.
-12(b), a smaller learning rate benefits SVFormer more, with warmup’s impact being comparatively small.
+15(a) and Fig.
+15(b), a smaller learning rate benefits SVFormer more, with warmup’s impact being comparatively small.
 This could be attributed to the model’s outcomes being closely tied to the total summed learning rate, which has weak connection with warmup steps (Kaplan et al., 2020).
@@
 Except for the 2M model, Llama-like models ranging from 82M to 468M, as well as models with the GPT2 architecture, exhibit similar critical points and final losses (see Fig.
-12(c) and Fig.
-12(d)).
+15(c) and Fig.
+15(d)).
 This suggests that the difference between SVFormer and the vanilla Transformer is not sensitive to model size and architecture.
@@
 We first observe the effects of sharing the first layer’s queries or keys across all layers in Fig.
-14, finding that this significantly impacts model performance, similar to the results in Fig.
+17, finding that this significantly impacts model performance, similar to the results in Fig.
 4.
 Additionally, sharing the first layer’s values in a multi-layer network may reduce the network’s “effective depth.” By updating the shared values using intermediate layers as “anchors,” we find that increasing the number of “anchors” improves performance, as shown in Fig.
-14.
+17.
 5 Conclusion
 In this paper, we propose the concept of attention concentration, a problem that arises from stacking multiple attention layers.
-From the perspective of cross-layer attention, we derive ResFormer, which adds a residual connection between the value vectors of the current layer and those of the first layer before the attention operation to alleviate attention concentration.
+From the perspective of cross-layer attention, we derive ResFormer, which adds a residual connection between the value states of the current layer and those of the first layer before the attention operation to alleviate attention concentration.
 Additionally, we introduce SVFormer, based on ResFormer, which reduces the $KV$ cache by nearly half.
@@
 These trends suggest that attention concentration indeed leads to the phenomenon of over-smoothing.
-Figure 15: (Left) The average token similarity of hidden states across layers in ResFormer vs.
-the vanilla Transformer.
-(Right) The average token similarity of hidden states across layers in Llama (8B) (Dubey et al., 2024) and Mistral (7B) (Jiang et al., 2023).
-A.2 Pre-train Dataset
+Figure 18: The average token similarity of hidden states across layers in Llama and Mistral.
+A.2 Attention Concentration Visualization
+We visualize the token importance, norms of value states and norms of hidden states for tokens at different position across layers.
+The results are averaged from 1,000 different sequences so that only the start token is the same and special across all sequences.
+Fig.
+19 (First row) demonstrates that the start token easily attracts massive attention despite lacking semantic information for Transformer and NeuTRENO.
+For Resformer, the importance of the start token is less than 10 times that of tokens at other positions, indicating that tokens carrying semantic information receive more attention.
+Moreover, both Transformer and NeuTRENO exhibit significant value-state drains (Guo et al., 2024b) and residual-state peaks (Guo et al., 2024a, Sun et al., 2024) on the start token at certain layers.
+In contrast, for Resformer, the value state norm of the start token is approximately 0.7 times that of other tokens, while the peak hidden state norm is about three times larger.
+Moreover, We conduct an entropy analysis on test examples of varying lengths.
+The results in Fig.20 indicate that Resformer is more effective in mitigating attention concentration when the sequences are shorter.
+Token
+importance
+Norms of
+value states
+Norms of
+hidden states
+(a) Transformer.
+(b) NeuTRENO.
+(c) Resformer.
+Figure 19: Visualization of the token importance, norms of value states and norms of hidden states.
+(a) Length: 20 tokens.
+(b) Length: 100 tokens.
+(c) Length: 500 tokens.
+(d) Length: 2,048 tokens.
+Figure 20: The average entropy of token importance across layers in ResFormer vs. the vanilla Transformer for sequence with different lengths.
+A.3 Ablation study of NeuTRENO
+NeuTRENO is sensitive to the choice of hyperparameter $\bm{\lambda}$ which is task-dependent.
+In the appendix of Nguyen et al.
+(2023), it is reported that $\lambda$ is set to 0.6 for image classification and segmentation tasks, and 0.4 for language modeling tasks.
+Fig.
+21 indicates that ${\bm{\lambda}}=0.4$ achieves the best results in our training dataset so that we choose ${\bm{\lambda}}=0.4$ for comparison.
+Besides, we empirically choose ${\bm{\lambda}}=0.2$ for NeuTRENO when combined with Resformer.
+Figure 21: Ablation study of different $\lambda$ for NeuTRENO.
+A.4 Pre-train Dataset
 Data source proportions Tokens
@@
 In this way, we sampled 20B tokens data from original 627B data based on the original data proportions shown in the Table 2.
-A.3 Training Details
+A.5 Training Details
 Max Sequence Length 512 2,048 8,192 32,000 64,000
@@
 Table 4: Training details for models with different size.
-A.4 Validation Loss On Slimpajama
+A.6 Validation Loss On Slimpajama
 Section 4.1.2 introduces to use relative training loss as a main evaluation matrix. Table 5 reports the validation loss for differnt model on the whole validation split of slimpajama.

### v3

@@
 To address this problem, we propose Transformer with residual value (ResFormer) which approximates cross-layer attention through adding a residual connection from the values of the the first layer to all subsequent layers.
-Based on this method, one variant is the Transformer with single layer value (SVFormer), where all layers share the same value embedding from first layer, reducing the $KV$ cache by nearly 50%.
-Comprehensive empirical evidence demonstrates that ResFormer mitigates attention concentration problem in deeper layers and enhances representation across most layers, outperforming the vanilla Transformer, DenseFormer, and NeuTRENO in training error as well as downstream tasks.
-Further visualization results suggest that Resformer alleviates attention sinks through avoiding value-state drains.
-SVFormer trains significantly faster than the vanilla Transformer and performs better than other methods like GQA and CLA, with performance influenced by sequence length and cumulative learning rate.
+Based on this method, one variant is the Transformer with single layer value (SVFormer), where all layers share the same value embedding from first layer.
+Comprehensive empirical evidence demonstrates ResFormer achieves equivalent validation loss with 10.4% fewer model parameters and 13.6% less training data compared to Transformer, while maintaining similar memory usage and computational cost.
+Besides, SVFormer reduces KV cache size by nearly half with only a small performance penalty and can be integrated with other KV-efficient methods, yielding further reductions in KV cache, with performance influenced by sequence length and cumulative learning rate.
+Further visualization results suggest that Resformer and SVFormer alleviate attention concentration in deeper layers through avoiding value-state drains and enhance representation across most layers.
 [^1] Code is available at https://github.com/Zcchill/Value-Residual-Learning.
+Figure 1: (Left) The average entropy of token importance and the average hidden-state similarity for a randomly initialized 468M model.
+(Middle) The average entropy of token importance across layers in Llama (8B) (Dubey et al., 2024) and Mistral (7B) (Jiang et al., 2023).
+(Right) The average entropy of token importance across layers in ResFormer vs.
+the vanilla Transformer, where token importance is derived from the attention matrix.
+Lower entropy indicates more focused attention on specific tokens.
+More details can be found in Eqn.
+11.
 1 Introduction
-The Transformer (Vaswani, 2017) model has become one of the leading architectures in recent years, excelling in both language modeling (Devlin, 2018, Lan, 2019, Brown, 2020) and computer vision tasks (Dosovitskiy, 2020).
+The Transformer (Vaswani et al., 2017) model has become one of the leading architectures in recent years, excelling in both language modeling (Devlin et al., 2019, Lan et al., 2020, Brown et al., 2020) and computer vision tasks (Dosovitskiy et al., 2021).
 The discovery of scaling laws (Hoffmann et al., 2022, Kaplan et al., 2020) has driven the pursuit of larger Transformer models by increasing network depth and width.
+(a) Token importance.
+(b) Norms of value states.
+(c) Norms of hidden states.
+Figure 2: The token importance (Xiao et al., 2024), value-state norms (Guo et al., 2024b), and hidden-state norms (Sun et al., 2024) of the first token across layers of 468M models.
+More Visualization results are available in Appendix A.2.
 Training large models presents significant challenges.
@@
 Recently, DenseFormer (Pagliardini et al., 2024) adopts the idea of DenseNet when training Transformer.
-Additionally, NeuTRENO (Nguyen et al., 2023) alleviates over-smoothing through incorporating the difference between the value states of the first layer and the current layer to the attention output.
-In this paper, we address the problem of multi-layer attention from a novel perspective.
-We introduce the phenomenon of attention concentration, which describes how a model’s attention increasingly focuses on fewer tokens as the network depth increases, as illustrated in Fig.
-1 (Right).
+Additionally, NeuTRENO (Nguyen et al., 2023) alleviates over-smoothing through incorporating the difference between the value vectors of the first layer and the current layer to the attention output.
+In this paper, we address the problem of multi-layer attention from another perspective.
+We introduce the phenomenon of attention concentration, which describes how a model’s attention increasingly focuses on fewer tokens.
 We quantify the degree of attention concentration using the entropy of the distribution of token importance, where lower entropy indicates a more pronounced concentration.
-In a deep model, the attention mechanism exhibits a pattern of “concentration - dispersion - concentration,” where the model could potentially lose useful information during the concentrated phase.
-See Fig.
-18 for analysis of over-smoothing.
-To alleviate the attention concentration from stacking multiple attention layers, an effective method is to use cross-layer attention on information from earlier layers.
-Given the high computational cost of cross-layer attention, we propose ResFormer as an approximation.
-ResFormer achieves a similar effect by applying a residual connection between the value states of the current layer and the first layer before the attention operation.
-Experimental results show that ResFormer mitigates the attention concentration effect in deeper layers, maintaining attention dispersion throughout the sequence, as shown in Fig.
-1 (Middle).
-Further visualization results show that Resformer effectively alleviates attention sinks (Xiao et al., 2023) through avoiding value-state drains (Guo et al., 2024b).
-During inference, deep networks require substantial $KV$ cache, severely impacting model deployment (Xiao et al., 2023).
+Unlike over-smoothing, which is inherent to model architecture, attention concentration emerges during training.
+Fig.
+1 (Left) shows that randomly initialized models exhibit over-smoothing but not attention concentration.
+Trained ViT models often focus on low-informative background areas (Darcet et al., 2024), while language models concentrate on low-semantic tokens (Sun et al., 2024), particularly the start token (attention sink (Xiao et al., 2024)).
+While previous studies analyzed single-layer attention patterns, our research reveals a “concentration - dispersion - concentration” pattern in deep models, as shown in Fig.
+1 (Middle), suggesting potential loss of information during concentrated phases.
+The analysis of over-smoothing is available in Appendix A.1.
+Mitigating attention concentration can lead to more interpretable attention maps and potentially improve downstream task performance (Darcet et al., 2024).
+This phenomenon typically emerges after the second or third network layer and is associated with value-state drains (decreased magnitude of value states) (Guo et al., 2024b), and hidden-state peaks (increased magnitude of hidden states) (Sun et al., 2024).
+Guo et al.
+(2024a) shows a mutual reinforcement mechanism exists between value-state drains and attention concentration.
+Recent studies have linked this to implicit biases during pretraining, with most existing solutions focusing on the use of additional tokens (registers) (Darcet et al., 2024) or additional keys and values (explicit attention bias) (Sun et al., 2024) to redirect this.
+Given that the first layer always shows no attention concentration, an effective method is to use cross-layer attention on information from this layer.
+However, due to computational costs, we propose ResFormer as an efficient alternative.
+ResFormer applies a residual connection between the value vectors of the current layer and the first layer before the attention operation.
+Unlike cross-layer attention, ResFormer indirectly mitigates attention concentration.
+It leverages the absence of value-state drains in the first layer by introducing a value residual connection.
+This alleviates value-state drains in deeper layers, thereby disrupting the mutual reinforcement between attention concentration and value-state drains, as shown in Fig.
+1 (Right) and Fig.
+2.
+During inference, deep networks require substantial $KV$ cache, severely impacting model deployment (Xiao et al., 2024).
 Existing $KV$-efficient methods often process keys and values simultaneously.
 Building on ResFormer, we decouple the value from the attention operation and propose a new kind of Transformer with single layer value (SVFormer).
-In SVFormer, the queries and keys of all layers share the value from the first layer.
-Consequently, SVFormer can save nearly half of the $KV$ cache.
-Additionally, SVFormer is orthogonal to the classical method GQA, and they can be used concurrently.
-Experimental results show that SVFormer trains significantly faster than the vanilla Transformer.
-Its performance mainly depends on factors like training sequence length and cumulative learning rate.
+In SVFormer, the queries and keys of all layers share the value from the first layer, and thus it can also alleviate attention concentration.
 We experiment on a 20B SlimPajama sub-sampled dataset, using settings similar to popular large language models (Wei et al., 2023, Dubey et al., 2024, Kaplan et al., 2020).
 We compare different models by their relative training curves against the vanilla Transformer.
-Results show that ResFormer outperforms the vanilla Transformer, DenseFormer, and NeuTRENO across all settings.
-Furthermore, SVFormer trains faster than the vanilla Transformer and performs better when the sequence length is longer.
-Figure 1: (Left) Illustration of the relative training loss (loss of target model - loss of vanilla Transformer) curve between different Transformer variants; model size is fixed to be 82M.
-(Middle) The average entropy of token importance across layers in ResFormer vs.
-the vanilla Transformer, where token importance is derived from the attention matrix.
-Lower entropy indicates more focused attention on specific tokens.
-More details can be found in Eqn.
-11.
-(Right) The average entropy of token importance across layers in Llama (8B) (Dubey et al., 2024) and Mistral (7B) (Jiang et al., 2023).
+Results show that ResFormer outperforms the vanilla Transformer, DenseFormer, and NeuTRENO.
+ResFormer achieves equivalent validation loss with 10.4% fewer model parameters and 13.6% less training data compared to Transformer, while maintaining similar memory usage and computational cost.
+Besides, SVFormer, while reducing the $KV$-cache by nearly half, requires a 12.2% increase in parameters to achieve the same validation loss as Transformer.
+And the performance of SVFormer is better when the training sequence length is longer.
+It further reduces the $KV$ cache when integrated with classical method GQA (Ainslie et al., 2023).
 (a) Transformer
@@
 Figure 3: Simplified illustration of the vanilla Transformer, NeuTRENO, DenseFormer, ResFormer, and SVFormer, with only three-layer structures and no operations other than attention.
-$A^{i}$, $V^{i}$, and $H^{i}$ denote the attention matrix, value states, and attention outputs at the $i$-th layer, respectively.
+$A^{i}$, $V^{i}$, and $H^{i}$ denote the attention matrix, value vectors, and attention outputs at the $i$-th layer, respectively.
 $\oplus$, $\ominus$, and $\otimes$ represent standard matrix addition, subtraction, and multiplication, respectively.
@@
 (2022) investigates over-smoothing from a graph perspective in BERT-based language modeling tasks.
-NeuTRENO (Nguyen et al., 2023) adds the difference between the value states of the first and current layers to each layer’s attention output and significantly alleviates the over-smoothing problem.
+NeuTRENO (Nguyen et al., 2023) adds the difference between the value vectors of the first and current layers to each layer’s attention output and significantly alleviates the over-smoothing problem.
 In contrast to these methods, ResFormer accesses and integrates information from previous layers prior to the attention operation, as illustrated in Fig.
-2.
+3.
 Moreover, it does not require the selection or tuning of additional hyperparameters.
@@
 Besides, several approaches have been proposed to enhance models’ ability to process long texts while reducing the reliance on $KV$ cache.
-Dai (2019) advocates segmenting long texts into smaller parts for attention computation.
-Building upon this, Munkhdalai et al.
-(2024) proposes using a fixed-size memory matrix for storing and retrieving historical information.
+Dai et al.
+(2019) advocates segmenting long texts into smaller parts for attention computation.
+Furthermore, Munkhdalai et al.
+(2024) uses a fixed-size memory matrix for storing and retrieving past information.
 Transformer-based methods can be categorized into three main groups.
@@
 Besides, CLA (Brandon et al., 2024) and LISA (Mu et al., 2024) respectively point out that we can reuse keys, values, or the attention matrix across layers to reduce redundancy between layers.
-While these methods typically process both key and value simultaneously, SVFormer is the first approach to decouple value from query and key.
-It shares values across all layers and is compatible with other methods like GQA.
+While these methods typically process both key and value simultaneously, SVFormer is the first approach to decouple value from query and key during attention computation.
+Moreover, it is compatible with other methods like GQA.
 3 Method
@@
 $\mathbf{U}_{n}$ $\approx\frac{1}{n}\mathbf{A}_{n}\sum_{i=1}^{n}{\mathbf{V}_{i}}.$ (5)
-Figure 3: Average token similarity between the outputs of different mapping methods and that of Eqn. 2.
-Figure 4: Ablation study on sharing keys or values in every two layers, with CLAttention denoting sharing both.
 where $\mathbf{A}_{n}$ denotes the original attention matrix for layer $n$.
@@
 3 projects the historical values into the current layer’s embedding space using the current layer’s attention as a weight matrix.
-For example, a naive approach would be to perform identity mapping, as described by
-$\mathbf{U}_{n}$ $=\mathbf{A}_{n}\mathbf{V}_{n}+\frac{1}{n-1}\sum_{i=1}^{n-1}{\mathbf{V}_{i}}.$ (6)
-To evaluate the approximation effect of replacing the $\mathbf{K}_{m}$ with $\mathbf{K}_{n}$, we randomly select 1,000 pre-training data samples.
-For each layer of a trained baseline model, assuming cross-layer attention defined by Eqn.
-2 is required for each layer with respect to the previous one, we calculate the cosine similarity between the outputs from Eqn.
+For example, a naive approach would be to perform identity mapping, as described by $\mathbf{U}_{n}$ $=\mathbf{A}_{n}\mathbf{V}_{n}+\frac{1}{n-1}\sum_{i=1}^{n-1}{\mathbf{V}_{i}}.$ (6) To evaluate the approximation effect of replacing the $\mathbf{K}_{m}$ with $\mathbf{K}_{n}$, we randomly select 1,000 pre-training data samples.
+For each layer of a trained baseline model, assuming cross-layer attention is required for each layer with respect to the previous one, we calculate the cosine similarity between the outputs from Eqn.
 2 and Eqn.
-3.
+4.
 We also calculate the cosine similarity between the outputs from Eqn.
@@
 Fig.
-4 shows that using current attention as a mapping matrix provides a better approximation for cross-layer attention.
+4 shows that our proposed method provides a good approximation for cross-layer attention.
+[Uncaptioned image] Figure 4: Average token similarity between layer outputs.
+Lines show similarity of outputs using current attention (Eqn.
+4) or identity attention (Eqn.
+6) compared to the one using cross-layer attention (Eqn.
+2).
 3.3 Transformer with Residual Value
@@
 Using our framework, the NeuTRENO can be defined as
-$\mathbf{U}_{n}$ $=\Big{(}\mathbf{A}_{n}-\bm{\lambda}\mathbf{J}\Big{)}{\mathbf{V}_{n}}+\bm{\lambda}{\mathbf{V}_{1}}.$ (8)
-where $\mathbf{J}$ denotes a matrix of ones and $\bm{\lambda}$ is a hyper-parameter.
-It can be found that the term of $\bm{\lambda}\mathbf{J}$ may have certain negative impact on the learning of original attention.
-If we ignore the attention output projections and the MLP layers, DenseFormer can also be modeled within our framework as
+$\mathbf{U}_{n}$ $=\Big{(}\mathbf{A}_{n}-\bm{\lambda}\mathbf{I}\Big{)}{\mathbf{V}_{n}}+\bm{\lambda}{\mathbf{V}_{1}}.$ (8)
+where $\mathbf{I}$ denotes the identity matrix and $\bm{\lambda}$ is a hyper-parameter.
+It can be found that the term of $\bm{\lambda}\mathbf{I}$ may have certain negative impact on the learning of original attention.
+If we ignore the attention output projection and the MLP layer, DenseFormer can also be modeled within our framework as
 $\mathbf{U}_{n}$ $=\sum_{i=1}^{n}{\bm{\alpha}_{i}\mathbf{A}_{i}\mathbf{V}_{i}}.$ (9)
 where $\{\bm{\alpha}_{i}\}_{i=1}^{n}$ is a set of hyper-parameters.
-DenseFormer uses attention matrix of previous layer as the weight matrix of projecting values but this is not aligned with the concept of the efficient cross layer attention shown in Eqn.
+DenseFormer uses attention matrix of previous layer as the weight matrix of projecting values but this is not aligned with the concept shown in Eqn.
 3.
 3.5 SVFormer: Single-layer Value For Half $KV$ Cache
-After ResFormer, a natural idea is whether we can remove the value states in each layer and have all layers share the value states from the first layer.
+After ResFormer, a natural idea is whether we can remove the value vectors in each layer and have all layers share the value vectors from the first layer.
 We call this method SVFormer.
@@
 $\mathbf{U}_{n}$ $=\mathbf{A}_{n}{\mathbf{V}_{1}}.$ (10)
-Compared to previous methods, SVFormer is the first method that decouple value states from attention.
-Its main advantage is that it only requires computing and storing the value states for the first layer, saving nearly half of the $KV$ cache during inference.
-Similar methods like CLA reduce $KV$ cache by sharing both of the key and value states every two layers, but the results in Fig.
-4 show that sharing values has less negative impact on performance compared with sharing keys.
+Compared to previous methods, SVFormer is the first method that decouple value vectors from attention.
+Its main advantage is that it only requires computing and storing the value vectors for the first layer, saving nearly half of the $KV$ cache during inference.
+Similar methods like CLA reduce $KV$ cache by sharing both of the key and value vectors every two layers.
+However, the results in Fig.
+5 show that sharing values has less negative impact compared with sharing keys.
+Figure 5: Ablation study of sharing different parts of attention every two layers.
 4 Pretrain Experiments
@@
 We adopt linear learning rate warmup for the first 1,200 steps with the initial learning rate and the peak learning rate to be 1e-7 and 6e-4 respectively.
-The cosine decay schedule gradually decays to 10% of the peak learning rate by the end of training (Brandon et al., 2024, Zhou et al., 2024, Wei et al., 2023).
-See Table 4 and Table 3 in Appendix for more details.
-All models are trained with 8 Nvidia A100 80G GPUs using mixed-precision training in FP16. We adopt deepspeed zero-2 optimizer and flash attention mechanism.
-Figure 5: (Left) The relative training curve between a 82M ResFormer and Transformer across different training sequence lengths.
-(Middle) Average training loss for the final 50 steps across different model sizes and the corresponding fitted curves.
-(Right) The relative training curve across different model size for a fixed 2,048 training sequence length.
+The cosine decay schedule gradually decays to 10% of the peak learning rate by the end of training (Zhou et al., 2024, Wei et al., 2023).
+The detailed hyperparameters for models of various sizes and different training sequence lengths in Appendix A.5.
+Moreover, All models are trained with 8 Nvidia A100 80G GPUs using mixed-precision training in FP16.
+We adopt deepspeed zero-2 optimizer and flash attention mechanism.
 4.1.2 Relative Training Loss Curve On SlimPajama
@@
 $\bm{e}$ $=-\sum_{i=1}^{l}\bm{a}_{i}^{\prime}\log{\bm{a}^{\prime}_{i}}.$ (11)
-where $\bm{a}_{i}^{\prime}={\bm{a}_{i}}/\Big{(}{\sum_{i=1}^{l}{\bm{a}_{i}}}\Big{)}$ for $i=1,2,\dots,l$ and the higher the entropy $\bm{e}$, the greater the degree of clustering in $\bm{a}$, i.e., attention matrix $\mathbf{A}$ is more likely to focus on several specific tokens.
+where $a_{i}^{{\prime}_{i}}={a_{i}}/\Big{(}{\sum_{i=1}^{l}{a_{i}}}\Big{)}$ for $i=1,2,\dots,l$ and the higher the entropy $\bm{e}$, the greater the degree of clustering in $\bm{a}$, i.e., attention matrix $\mathbf{A}$ is more likely to focus on several specific tokens.
 4.1.4 Spectral Decomposition For Analyzing Representations
@@
 Here we use spectral decomposition to analyze the feature space of value $\bm{v}$ of one layer as following:
-$\frac{1}{l}\sum_{i=1}^{l}\bm{v}_{i}\bm{v}_{i}^{T}$ $=\sum_{i=j}^{d}\bm{w}_{j}\bm{\lambda}_{j}\bm{w}_{j}^{T}.$ (12)
-where $\bm{w}_{j}$ is the $j$-th eigenvector with eigenvalue $\bm{\lambda}_{j}$ for $i=1,2,\dots,d$ and $d$ is the dimensionality of the value’s feature space.
+$\frac{1}{l}\sum_{i=1}^{l}\bm{v}_{i}\bm{v}_{i}^{T}$ $=\sum_{j=1}^{d}\bm{w}_{j}\bm{\lambda}_{j}\bm{w}_{j}^{T}.\vspace{-3mm}$ (12)
+where $\bm{w}_{j}$ is the $j$-th eigenvector with eigenvalue $\bm{\lambda}_{j}$ for $j=1,2,\dots,d$ and $d$ is the dimensionality of the value’s feature space.
+Figure 6: (Left) Validation loss as model size scales from 82M to 468M parameters.
+(Right) Validation loss for the 468M parameter model evaluated every 2B tokens.
+ResFormer achieves approximately 10.4%-13.6% reduction in both model parameters and training data.
+Figure 7: (Left) The relative training curve between a 82M ResFormer and Transformer across different training sequence lengths.
+(Right) Relative training loss of various Transformer variants compared to the vanilla Transformer model, with model size fixed at 82M parameters.
 4.2 ResFormer vs. vanilla Transformer
@@
 In Fig.
-5, ResFormer consistently outperforms vanilla Transformer throughout training for different models sizes.
+7 (Left), ResFormer consistently outperforms vanilla Transformer throughout training across different training sequence lengths.
 Additionally, the results in Fig.
-1 (Left) illustrate that ResFormer outperforms DenseFormer and NeuTRENO.
+7 (Right) illustrate that ResFormer outperforms DenseFormer and NeuTRENO.
 Furthermore, integrating ResFormer with NeuTRENO leads to additional performance improvements.
+We also analyzed how ResFormer and Transformer scale at model size and data size.
+ResFormer and Transformer are trained on similar experiment setting.
+On the one hand, we trained model with 82M, 180M, 320M and 468M parameters on 20B training tokens and evaluated them on a separate validation set.
+As shown in Fig.7 (Left), ResFormer achieves equivalent validation loss to the Transformer while utilizing 10.4% fewer model parameters.
+On the other hand, we evaluated the 468M models every 2B tokens and ResFormer needs 13.6% fewer training tokens to achieve the same validation loss as Transformer.
+The validation loss for these models is available in Appendix A.6.
 We further test the variant of ResFormer defined as $\mathbf{U}_{n}=\mathbf{A}_{n}({\mathbf{V}_{n}}+{\bm{\lambda}}{\mathbf{V}_{1}})$.
@@
 Regardless of the value of $\bm{\lambda}$, ResFormer consistently outperforms Transformers.
-It suggests that the success of Resformer lies in the use of $\mathbf{V}_{1}$ and the transformation through $\mathbf{A}_{n}$.
+It suggests that the success of ResFormer lies in the use of $\mathbf{V}_{1}$ and the mapping by $\mathbf{A}_{n}$.
 The ablation study of different hyperparameters $\bm{\lambda}$ for NeuTRENO, as defined in Equation 8, can be found in the Appendix A.3.
-[Uncaptioned image] Figure 6: Ablation study of different $\bm{\lambda}$ for Resformer.
+Figure 8: Ablation study of different $\bm{\lambda}$ for ResFormer.
 Figure 9: Ablation study of adding residual connection to queries or keys.
@@
 4.3 Ablation study of residual connection
-(a) Token importance.
-(b) Norms of value states.
-(c) Norms of hidden states.
-Figure 10: The token importance, value-state norms, and hidden-state norms of the first token in each sequence across layers in ResFormer vs. the vanilla Transformer.
-(a) Transformer.
-(b) NeuTRENO.
-(c) Resformer.
-Figure 11: The distribution of token importance for different models at different layers.
 In Eqn.
@@
 The results, shown in Fig.
-9, indicate that only residual connections for values yield positive effects.
-One possible explanation is that attention mechanisms are sensitive to perturbations, and modifying queries or keys can significantly impact attention matrix.
+11, indicate that only residual connections for values yield positive effects.
+One possible explanation is that attention mechanisms are sensitive to perturbations, and modifying queries or keys significantly impacts it.
 Moreover, we compare with the models based on Eqn.
 2 and Eqn.
-6.
-From the cross layer attention perspective, the former uses a gold attention matrix in Eqn.
-2 to map historical values, while the latter uses an all-ones matrix in Eqn.
 6.
@@
 4, showing that identity mapping causes significant perturbations, leading to poor performance.
-Interestingly, using current attention as the mapping matrix results in an even lower final loss than using gold attention.
+Interestingly, ResFormer achieves an even lower final loss than ResFormer.
+It suggests that ResFormer’s impact on the attention optimization is better by mitigating value-state drains.
 When determining the mapping method and target value, it is crucial to consider which historical layers’ values should be included in the residual connection.
@@
 The tasks include Hellaswag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2019), ARC-Easy and ARC-Challenge (Clark et al., 2018) and PIQA (Bisk et al., 2020).
-The results in Table 1 show that ResFormer (82M) achieves an average accuracy improvement of nearly 3% compared to the vanilla Transformer (82M).
-Model Max Length HellaSwag Obqa WinoGrande ARC-c ARC-e PIQA Avg.
+The results in Table 1 show that ResFormer achieved an average accuracy improvement of nearly 3% compared to the vanilla Transformer.
+Model Max Length HellaSwag Obqa WinoGrande ARC-c ARC-e PIQA Avg
 Transformer 2,048 0.263 0.142 0.492 0.199 0.331 0.572 0.333
@@
 Table 1: Zero-shot accuracy on commonsense reasoning tasks.
-Figure 12: Left: Distribution of eigenvalues for the value states in the first layer of ResFormer and Transformer. Right: Maximum eigenvalue for each layer of ResFormer and Transformer.
+Figure 12: Left: Distribution of eigenvalues for the value vectors in the first layer of ResFormer and Transformer. Right: Maximum eigenvalue for each layer of ResFormer and Transformer.
 4.5 Visualization of ResFormer
@@
 12 indicate that the value states generated by most layers of the ResFormer exhibit stronger representational capacity compared to those of the vanilla Transformer.
-Besides, we analyze the attention concentration effects mentioned in Section 4.1.3 using the same batch of test data.
+We also analyze the attention concentration effects mentioned in Section 4.1.3 using the same batch of test data.
 Fig.
-1 (Middle) illustrates that the clustering effect of attention increases significantly with the number of layers for the vanilla Transformer, whereas the clustering effect is relatively less pronounced for the ResFormer.
+1 (Right) illustrates that the clustering effect of attention increases significantly with the number of layers for the vanilla Transformer, whereas the clustering effect is relatively less pronounced for the ResFormer.
 We further visualize the attention weights, value-state norms $\|\bm{v}\|_{2}$, and hidden-state norms $\|\bm{h}\|_{2}$ of tokens at different layers and positions, with detailed results in Appendix A.2.
 Given that attention clustering often occurs on the first token, we primarily show its results in Fig.
-10.
-The results indicate that using ResFormer significantly mitigates attention sinks (Xiao et al., 2023), value-state drains (Guo et al., 2024b) and residual-state peaks (Sun et al., 2024).
+2.
+The results indicate that using ResFormer significantly mitigates attention sinks (Xiao et al., 2024), value-state drains (Guo et al., 2024b) and residual-state peaks (Sun et al., 2024).
 Guo et al.
 (2024a) attributes these phenomena to the mutual reinforcement mechanism of model and we suggest that the value shortcut disrupts this mechanism by alleviating value-state drains.
-Specifically, for tokens lacking semantic information, such as the first token, a large value state magnitude can adversely affect the prediction of future tokens if it is overly attended to, resulting in higher loss.
+Specifically, for tokens lacking semantic information like start tokens, a large value state magnitude can adversely affect the prediction of subsequent tokens if they are overly attended to.
+When there is no value-state drains, models will reduce attention clustering to these tokens to minimize loss.
 4.6 SVFormer vs. GQA
-In the Fig.
-13, at a training sequence length of 64,000, SVFormer demonstrates lower final loss compared to existing KV-efficient methods such as CLA and GQA.
-Moreover, it can be used concurrently with GQA to enhance KV efficiency further.
-However, we observed that with a training sequence length of 2,048, SVFormer underperforms compared to GQA.
-The results indicate that sequence length significantly affects SVFormer’s performance.
-Thus, we conducted more comprehensive experiments on sequence length.
-Results in Fig.
-14 (Left) demonstrate that SVFormer will always be gradually surpassed by vanilla attention during training while its training speed is much faster than vanilla attention.
-However, as the training sequence length increases, the SVFormer model performs better.
-In this way, we focus on the critical point, defined as the number of training steps exceeded.
-Fig.
-14 (Right) illustrates that the relationship between the critical point and sequence length exhibits an exponential trend.
-We argue that it’s due to the challenge deep models face in fully optimizing the increasingly larger first-layer value matrix as the training sequence length grows.
 Figure 13: The relative training loss for SVFormer and other $KV$ efficient model compared with vanilla attention.
@@
 Right: Analysis of critical point, and we predict it for length 64,000 using linear regression with the last 1,000 data points.
+In the Fig.
+14, at a training sequence length of 64,000, SVFormer demonstrates lower final loss compared to existing $KV$-efficient methods such as CLA and GQA.
+Moreover, it can be used concurrently with GQA to enhance $KV$ efficiency further.
+However, we observed that with a training sequence length of 2,048, SVFormer underperforms compared to GQA.
+The results indicate that sequence length significantly affects SVFormer’s performance.
+Thus, we conducted more comprehensive experiments on sequence length.
+Results in Fig.
+14 (Left) demonstrate that SVFormer will always be gradually surpassed by vanilla attention during training while its training speed is faster than vanilla Transformer at the early stage.
+However, as the training sequence length increases, the SVFormer model performs better.
+In this way, we focus on the critical point, defined as the number of training steps exceeded.
+Fig.
+14 (Right) illustrates that the relationship between the critical point and sequence length exhibits an exponential trend.
+We argue that it’s due to the challenge deep models face in fully optimizing the increasingly larger first-layer value matrix as the training sequence length grows.
 4.7 Other Factors Influencing SVFormer
@@
 Moreover, larger models often require smaller learning rates to ensure training stability, making them more suitable for using SVFormer.
-Except for the 2M model, Llama-like models ranging from 82M to 468M, as well as models with the GPT2 architecture, exhibit similar critical points and final losses (see Fig.
-15(c) and Fig.
-15(d)).
-This suggests that the difference between SVFormer and the vanilla Transformer is not sensitive to model size and architecture.
+Llama-like models and GPT2-like models exhibit similar critical points and final losses (see Fig.
+15(c)).
+This suggests that the difference between SVFormer and the vanilla Transformer is not sensitive to architecture.
+Compared with Transformer, SVFormer requires a 12.2% increase in parameters to achieve the same validation loss while reducing the $KV$-cache by nearly half.
 (a) Learning Rate.
 (b) Warmup Steps.
-(c) Model Size.
-(d) Architecture.
-Figure 15: The relative training loss for SVFormer under different hyper-parameter setting.
+(c) Architecture.
+(d) Model Size.
+Figure 15: The relative training loss for SVFormer under different hyper-parameter setting and the validation loss as model size scales from 82M to 468M parameters.
 4.8 Ablation study of SVFormer
@@
 17, finding that this significantly impacts model performance, similar to the results in Fig.
-4.
+5.
 Additionally, sharing the first layer’s values in a multi-layer network may reduce the network’s “effective depth.” By updating the shared values using intermediate layers as “anchors,” we find that increasing the number of “anchors” improves performance, as shown in Fig.
@@
 In this paper, we propose the concept of attention concentration, a problem that arises from stacking multiple attention layers.
-From the perspective of cross-layer attention, we derive ResFormer, which adds a residual connection between the value states of the current layer and those of the first layer before the attention operation to alleviate attention concentration.
+From the perspective of cross-layer attention, we derive ResFormer, which adds a residual connection between the value vectors of the current layer and those of the first layer before the attention operation to alleviate attention concentration.
 Additionally, we introduce SVFormer, based on ResFormer, which reduces the $KV$ cache by nearly half.
 We conducted comprehensive experiments on the language modeling task to validate the advantages of these two Transformer variants in different scenarios.
-Acknowledgement
-This work was supported by the Research Center for Industries of the Future at Westlake University (Grant No. WU2023C017) and the Key Research.
 Ethics Statement
@@
 Reproducibility Statement
-We have detailed the complete experiments setup such as batch size, optimizer and learning rates in Section 4.1.1.
-Besides, we have released the source codes.
-These resources should be sufficient to reproduce the results of the paper.
+We have detailed the complete experiments setup such as batch size, optimizer, learning rates in Section 4.1.1.
+Besides, we will release source codes once our paper is made public.
+These resources should be sufficient to reproduce results of the paper.
 Appendix A Appendix
@@
 In the case of Llama and Mistral, the average token similarity demonstrates an “M”-shaped pattern with increasing network depth, while entropy follows a “W”-shaped pattern at corresponding positions.
-These trends suggest that attention concentration indeed leads to the phenomenon of over-smoothing.
+These trends indicate a strong correlation between attention concentration and over-smoothing.
 Figure 18: The average token similarity of hidden states across layers in Llama and Mistral.
 A.2 Attention Concentration Visualization
+Transformer
+NeuTRENO
+ResFormer
+SVFormer
+(a) Token importance.
+(b) Value-state norms.
+(c) Hidden-state drains.
+Figure 19: Visualization of token importance, value state norms, and hidden state norms across different token positions and layers in 468M models.
+(a) Entropy.
+(b) Token importance.
+(c) Value-state norms.
+(d) Hidden-state norms.
+Figure 20: Token analysis in 82M models. (a) Importance entropy of sequences; (b) Token importance, (c) value-state norms, and (d) hidden-state norms of the first token across layers.
+(a) Transformer.
+(b) NeuTRENO.
+(c) ResFormer.
+(d) SVFormer.
+Figure 21: The distribution of token importance for different models at different layers.
 We visualize the token importance, norms of value states and norms of hidden states for tokens at different position across layers.
@@
 Fig.
-19 (First row) demonstrates that the start token easily attracts massive attention despite lacking semantic information for Transformer and NeuTRENO.
+19 (First column) demonstrates that the start token easily attracts massive attention despite lacking semantic information for Transformer and NeuTRENO.
 For ResFormer, the importance of the start token is less than 10 times that of tokens at other positions, indicating that tokens carrying semantic information receive more attention.
 Moreover, both Transformer and NeuTRENO exhibit significant value-state drains (Guo et al., 2024b) and residual-state peaks (Guo et al., 2024a, Sun et al., 2024) on the start token at certain layers.
-In contrast, for Resformer, the value state norm of the start token is approximately 0.7 times that of other tokens, while the peak hidden state norm is about three times larger.
-Moreover, We conduct an entropy analysis on test examples of varying lengths.
-The results in Fig.20 indicate that Resformer is more effective in mitigating attention concentration when the sequences are shorter.
-Token
-importance
-Norms of
-value states
-Norms of
-hidden states
-(a) Transformer.
-(b) NeuTRENO.
-(c) Resformer.
-Figure 19: Visualization of the token importance, norms of value states and norms of hidden states.
-(a) Length: 20 tokens.
-(b) Length: 100 tokens.
-(c) Length: 500 tokens.
-(d) Length: 2,048 tokens.
-Figure 20: The average entropy of token importance across layers in ResFormer vs. the vanilla Transformer for sequence with different lengths.
+In contrast, for ResFormer, the value state norm of the start token exceeds half the magnitude of other tokens, while the peak hidden state norm is less than triple the average.
+Fig.21 further illustrates the distribution of token importance, where TOP-$i$ represents the $i$-th largest token importance within a sequence.
+Compared to Transformer and NeuTRENO, ResFormer and SVFormer exhibit a more uniform distribution of token importance.
+Similar to Fig.2, we conducted experiments on 82M models, with results shown in Fig.20.
+We also illustrate the attention pattern of cross-layer attention introduced in Eqn.
+2.
+The results demonstrate that while cross-layer attention successfully mitigates the problem of attention concentration, it still exhibits value-state drains.
 A.3 Ablation study of NeuTRENO

### v4

@@
-Value Residual Learning For Alleviating Attention Concentration In Transformers
+Value Residual Learning
 Abstract
-Transformers can capture long-range dependencies using self-attention, allowing tokens to attend to all others directly.
-However, stacking multiple attention layers leads to attention concentration.
-One natural way to address this issue is to use cross-layer attention, allowing information from earlier layers to be directly accessible to later layers.
-However, this approach is computationally expensive.
-To address this problem, we propose Transformer with residual value (ResFormer) which approximates cross-layer attention through adding a residual connection from the values of the the first layer to all subsequent layers.
-Based on this method, one variant is the Transformer with single layer value (SVFormer), where all layers share the same value embedding from first layer.
-Comprehensive empirical evidence demonstrates ResFormer achieves equivalent validation loss with 10.4% fewer model parameters and 13.6% less training data compared to Transformer, while maintaining similar memory usage and computational cost.
+While Transformer models have achieved remarkable success in various domains, the effectiveness of information propagation through deep networks remains a critical challenge.
+Standard hidden state residuals often fail to adequately preserve initial token-level information in deeper layers.
+This paper introduces ResFormer, a novel architecture that enhances information flow by incorporating value residual connections in addition to hidden state residuals.
+And a variant is the SVFormer, where all layers share the first layer’s value embedding.
+Comprehensive empirical evidence demonstrates ResFormer achieves equivalent validation loss with 13.3% fewer model parameters and 15.4% less training data compared to Transformer, while maintaining similar memory usage and computational cost.
 Besides, SVFormer reduces KV cache size by nearly half with only a small performance penalty and can be integrated with other KV-efficient methods, yielding further reductions in KV cache, with performance influenced by sequence length and cumulative learning rate.
-Further visualization results suggest that Resformer and SVFormer alleviate attention concentration in deeper layers through avoiding value-state drains and enhance representation across most layers.
-[^1] Code is available at https://github.com/Zcchill/Value-Residual-Learning.
-Figure 1: (Left) The average entropy of token importance and the average hidden-state similarity for a randomly initialized 468M model.
-(Middle) The average entropy of token importance across layers in Llama (8B) (Dubey et al., 2024) and Mistral (7B) (Jiang et al., 2023).
-(Right) The average entropy of token importance across layers in ResFormer vs.
-the vanilla Transformer, where token importance is derived from the attention matrix.
-Lower entropy indicates more focused attention on specific tokens.
-More details can be found in Eqn.
-11.
+Value Residual Learning
+Zhanchao Zhou1,2   Tianyi Wu3†thanks: Equal Contribution; †Work done during internship at Westlake University; $\diamond$ Corresponding author.
+Zhiyun Jiang4†∗   Fares Obeid5   Zhenzhong Lan2⋄ 1Zhejiang University      2Westlake University 3University of Electronic Science and Technology of China 4China University of Mining and Technology     5Imperial College London
 1 Introduction
-The Transformer (Vaswani et al., 2017) model has become one of the leading architectures in recent years, excelling in both language modeling (Devlin et al., 2019, Lan et al., 2020, Brown et al., 2020) and computer vision tasks (Dosovitskiy et al., 2021).
-The discovery of scaling laws (Hoffmann et al., 2022, Kaplan et al., 2020) has driven the pursuit of larger Transformer models by increasing network depth and width.
-(a) Token importance.
-(b) Norms of value states.
-(c) Norms of hidden states.
-Figure 2: The token importance (Xiao et al., 2024), value-state norms (Guo et al., 2024b), and hidden-state norms (Sun et al., 2024) of the first token across layers of 468M models.
-More Visualization results are available in Appendix A.2.
-Training large models presents significant challenges.
-Balancing the depth and width of a Transformer model within a fixed parameter budget is particularly difficult.
-While research indicates that deeper models generalize more compositionally than shallower ones (Petty et al., 2024), the training and deployment of deep models remain problematic.
-Although Transformers use residual connections (He et al., 2016) to address the vanishing gradient issue, training very deep Transformers is still challenging.
-For example, a 32-layer Vision Transformer (ViT) may perform worse than a 24-layer one (Zhou et al., 2021).
-This is mainly due to the smoothing mechanism of attention (Shi et al., 2022), which can lead to an over-smoothing effect (Nguyen et al., 2023) where the token representations become the same as the model’s depth increases.
-Existing solutions to alleviate the over-smoothing problem in Transformer include adding extra regularizers (Nguyen et al., 2023, Shi et al., 2022) and optimizing the information flow within the model (Pagliardini et al., 2024).
-During the era of convolutional neural network architectures, Stochastic Depth (Huang et al., 2016) reduces the likelihood of over-smoothing by randomly dropping layers during training and DenseNet (Huang et al., 2017) mitigates the impact of over-smoothing by allowing each layer to directly access the hidden states of all preceding layers.
-Recently, DenseFormer (Pagliardini et al., 2024) adopts the idea of DenseNet when training Transformer.
-Additionally, NeuTRENO (Nguyen et al., 2023) alleviates over-smoothing through incorporating the difference between the value vectors of the first layer and the current layer to the attention output.
-In this paper, we address the problem of multi-layer attention from another perspective.
-We introduce the phenomenon of attention concentration, which describes how a model’s attention increasingly focuses on fewer tokens.
-We quantify the degree of attention concentration using the entropy of the distribution of token importance, where lower entropy indicates a more pronounced concentration.
-Unlike over-smoothing, which is inherent to model architecture, attention concentration emerges during training.
-Fig.
-1 (Left) shows that randomly initialized models exhibit over-smoothing but not attention concentration.
-Trained ViT models often focus on low-informative background areas (Darcet et al., 2024), while language models concentrate on low-semantic tokens (Sun et al., 2024), particularly the start token (attention sink (Xiao et al., 2024)).
-While previous studies analyzed single-layer attention patterns, our research reveals a “concentration - dispersion - concentration” pattern in deep models, as shown in Fig.
-1 (Middle), suggesting potential loss of information during concentrated phases.
-The analysis of over-smoothing is available in Appendix A.1.
-Mitigating attention concentration can lead to more interpretable attention maps and potentially improve downstream task performance (Darcet et al., 2024).
-This phenomenon typically emerges after the second or third network layer and is associated with value-state drains (decreased magnitude of value states) (Guo et al., 2024b), and hidden-state peaks (increased magnitude of hidden states) (Sun et al., 2024).
-Guo et al.
-(2024a) shows a mutual reinforcement mechanism exists between value-state drains and attention concentration.
-Recent studies have linked this to implicit biases during pretraining, with most existing solutions focusing on the use of additional tokens (registers) (Darcet et al., 2024) or additional keys and values (explicit attention bias) (Sun et al., 2024) to redirect this.
-Given that the first layer always shows no attention concentration, an effective method is to use cross-layer attention on information from this layer.
-However, due to computational costs, we propose ResFormer as an efficient alternative.
-ResFormer applies a residual connection between the value vectors of the current layer and the first layer before the attention operation.
-Unlike cross-layer attention, ResFormer indirectly mitigates attention concentration.
-It leverages the absence of value-state drains in the first layer by introducing a value residual connection.
-This alleviates value-state drains in deeper layers, thereby disrupting the mutual reinforcement between attention concentration and value-state drains, as shown in Fig.
-1 (Right) and Fig.
-2.
-During inference, deep networks require substantial $KV$ cache, severely impacting model deployment (Xiao et al., 2024).
-Existing $KV$-efficient methods often process keys and values simultaneously.
-Building on ResFormer, we decouple the value from the attention operation and propose a new kind of Transformer with single layer value (SVFormer).
-In SVFormer, the queries and keys of all layers share the value from the first layer, and thus it can also alleviate attention concentration.
-We experiment on a 20B SlimPajama sub-sampled dataset, using settings similar to popular large language models (Wei et al., 2023, Dubey et al., 2024, Kaplan et al., 2020).
-We compare different models by their relative training curves against the vanilla Transformer.
-Results show that ResFormer outperforms the vanilla Transformer, DenseFormer, and NeuTRENO.
-ResFormer achieves equivalent validation loss with 10.4% fewer model parameters and 13.6% less training data compared to Transformer, while maintaining similar memory usage and computational cost.
-Besides, SVFormer, while reducing the $KV$-cache by nearly half, requires a 12.2% increase in parameters to achieve the same validation loss as Transformer.
-And the performance of SVFormer is better when the training sequence length is longer.
-It further reduces the $KV$ cache when integrated with classical method GQA (Ainslie et al., 2023).
+The Transformer (Vaswani et al., 2017) model has become one of the leading architectures in recent years, excelling in both language modeling (Devlin et al., 2019; Lan et al., 2020; Brown et al., 2020) and computer vision tasks (Dosovitskiy et al., 2021).
+Among its variants, decoder-only architectures have become the most prominent (Kaplan et al., 2020; Dubey et al., 2024).
+The discovery of scaling laws (Hoffmann et al., 2022; Kaplan et al., 2020) has driven the pursuit of larger Transformer models by increasing network depth and width.
+In a standard decoder-only transformer, initial token embeddings contain localized information, which rapidly e volves into abstract semantic features through early attention layers (Sun et al., 2024b; Clark et al., 2019).
+As Transformer deepens, a critical question arises: How effectively is the initial information propagated to deeper layers?
+One common answer is that residual connections of hidden states ensure access to initial information throughout the network.
+However, some studies (Zhou et al., 2021; Shi et al., 2022) have identified that the smoothing effect of attention mechanisms leads to over-smoothing, where token representations become increasingly similar as the network deepens.
+This indicates that in deeper layers, sequence-level features become dominant, while token-level features are diluted.
+DenseFormer (Pagliardini et al., 2024) applied the idea of learnable dense connections from DenseNet (Huang et al., 2016) to transformers, and the learned connection coefficients shows that deeper layers indeed require larger attention to initial embeddings.
+Given the low similarity between initial token embeddings and deeper hidden states (Sun et al., 2024b), their directly summation may significantly impact the modeling of attention distribution for abstract semantic information in later layers.
+NeuTRENO (Nguyen et al., 2023) alleviates over-smoothing from the view of regularizers by considering the difference between value vectors of the first and current layers.
 (a) Transformer
@@
 Figure 1: Simplified illustration of the vanilla Transformer, NeuTRENO, DenseFormer, ResFormer, and SVFormer, with only three-layer structures and no operations other than attention.
-$A^{i}$, $V^{i}$, and $H^{i}$ denote the attention matrix, value vectors, and attention outputs at the $i$-th layer, respectively.
+$\mathbf{A}^{i}$, $\mathbf{V}^{i}$, and $\mathbf{H}^{i}$ denote the attention matrix, value vectors, and attention outputs at the $i$-th layer, respectively.
 $\oplus$, $\ominus$, and $\otimes$ represent standard matrix addition, subtraction, and multiplication, respectively.
+In this paper, we propose ResFormer, which enhances the propagation of initial local information by introducing value residual connections in addition to the standard hidden residual connections.
+Specifically, ResFormer applies a residual connection between the value vectors of the current layer and the first layer before the attention operation.
+In other words, both value states share the existing attention matrix of the current layer.
+The value states of the first attention layer and the preceding hidden states differ only by a linear transformation along the channel dimension, both representing token-level raw information.
+We hypothesize that introducing residual connections for values has a less impact on modeling attention distributions for sequence-level semantic information in higher layers and complements the original hidden state residual.
+Fig.
+1 illustrates a comparison of the extra skip connections introduced by different models.
+During inference, deep networks require substantial $KV$ cache, severely impacting model deployment (Xiao et al., 2024).
+Existing $KV$-efficient methods often process keys and values simultaneously.
+Building on ResFormer, we decouple the value from the attention operation and propose a new kind of Transformer (SVFormer) where all layers share a single value states.
+We experiment on a 20B SlimPajama sub-sampled dataset, using settings similar to popular large language models (Wei et al., 2023; Dubey et al., 2024; Kaplan et al., 2020).
+We compare different models based on the valid loss against the vanilla Transformer.
+Results show that ResFormer outperforms the vanilla Transformer, DenseFormer, and NeuTRENO.
+ResFormer achieves equivalent validation loss with 13.3% fewer model parameters and 15.4% less training data compared to Transformer, while maintaining similar memory usage and computational cost.
+Besides, SVFormer, while reducing the $KV$-cache by nearly half, requires a 12.2% increase in parameters to achieve the same validation loss as Transformer.
+And SVFormer performs better when the training sequence length is longer.
+It further reduces the $KV$ cache when integrated with GQA (Ainslie et al., 2023).
 2 Related Work
-2.1 Shortcut Connections For Better Information Flow
+2.1 Shortcut Connections
 Deep learning models often consist of multiple layers, posing a challenge to minimize information loss during transmission.
@@
 These two methods further enhance the information flow after ResNet.
-Related research indicates that for advanced Transformer architectures, although increasing depth continues to yield performance improvements in language modeling tasks, the gains become less significant with further increases (Petty et al., 2024).
+Related research indicates that, although increasing depth continues to yield performance improvements in language modeling tasks, the gains become less significant with further increases (Petty et al., 2024).
 Furthermore, Zhou et al.
 (2021) illustrates that a 32-layer ViT underperforms a 24-layer ViT.
-Depth-Wise Attention (ElNokrashy et al., 2024) allows each query to access the key and value at the same position from previous layers through an attention-like mechanism before the output layer.
 DenseFormer (Pagliardini et al., 2024) integrates weighted fusion of outputs from all preceding layers after each layer.
@@
 NeuTRENO (Nguyen et al., 2023) adds the difference between the value vectors of the first and current layers to each layer’s attention output and significantly alleviates the over-smoothing problem.
-In contrast to these methods, ResFormer accesses and integrates information from previous layers prior to the attention operation, as illustrated in Fig.
-3.
-Moreover, it does not require the selection or tuning of additional hyperparameters.
 2.2 $KV$ cache compressing
-The $KV$ cache is a key factor limiting the efficiency of long-text model inference.
-Research in this area can be broadly classified into Transformer-based methods, which target redundant information in Transformer models, and non-Transformer methods, which mainly addresses the quadratic time complexity of attention with respect to sequence length.
-For non-Transformer methods, Mamba (Gu & Dao, 2023) and RWKV (Peng et al., 2023) are two popular works.
-They replace the original softmax-based attention with SSM (Gu et al., 2021) and AFT (Zhai et al., 2021) mechanisms, respectively.
-Besides, several approaches have been proposed to enhance models’ ability to process long texts while reducing the reliance on $KV$ cache.
-Dai et al.
-(2019) advocates segmenting long texts into smaller parts for attention computation.
-Furthermore, Munkhdalai et al.
-(2024) uses a fixed-size memory matrix for storing and retrieving past information.
-Transformer-based methods can be categorized into three main groups.
-The first group consists of post-training methods like SnapKV (Li et al., 2024) and ThinK (Xu et al., 2024), which compress $KV$ cache during inference based on attention matrices at token or hidden dimension levels.
-The second group focuses on quantization and adopts low-precision $KV$ cache quantization rather than completely eliminating them (Hooper et al., 2024).
-The third group aims to maximize the efficiency of attention-based models via parameter or activation value sharing.
+The KV cache significantly impacts the efficiency of long-text model inference, attracting extensive research.
+One category of Transformer-based methods addresses this by employing parameter or activation value sharing techniques.
 The most representative works include Multi-Query Attention (Shazeer, 2019) and Grouped-Query Attention (Ainslie et al., 2023) which suggest to share key and value across a group of queries.
-MLKV (Zuhri et al., 2024) further suggest to share keys and values for queries across layers and MLA (Liu et al., 2024) introduces low-rank projection when processing keys and values.
 Besides, CLA (Brandon et al., 2024) and LISA (Mu et al., 2024) respectively point out that we can reuse keys, values, or the attention matrix across layers to reduce redundancy between layers.
-While these methods typically process both key and value simultaneously, SVFormer is the first approach to decouple value from query and key during attention computation.
-Moreover, it is compatible with other methods like GQA.
+While these methods typically process both key and value simultaneously, SVFormer is the first approach to decouple value from query and key during attention.
 3 Method
-3.1 Motivation: Information Transfer Via Cross Layer Attention
+3.1 Notations
 Let $\mathbf{H}_{n}\in\mathbb{R}^{l\times d}$ be the input hidden state of the $n$-th layer, where $l$ denotes the sequence length and $d$ is the dimension size.
-In standard attention, the hidden state $\mathbf{H}_{n}$ will be firstly projected into $\mathbf{Q}_{n},\mathbf{K}_{n},\mathbf{V}_{n}\in\mathbb{R}^{l\times d}$ through three linear projections $\mathbf{W^{\mathbf{Q}}},\mathbf{W^{\mathbf{K}}},\mathbf{W^{\mathbf{V}}}\in\mathbb{R}^{d\times d}$ respectively.
-For simplicity, we introduce dot-product attention of layer $n$ as
-$\operatorname{Attention}(\mathbf{Q}_{n},\mathbf{K}_{n},\mathbf{V}_{n})$ $=\operatorname{Softmax}(\frac{\mathbf{Q}_{n}\mathbf{K}_{n}^{T}}{\sqrt{d}})\mathbf{V}_{n}.$ (1)
-An ideal way to incorporate previous layers’ information is cross layer attention.
-The attention mechanism naturally extracts relevant information from previous layers.
-If these layers contain low-quality information, the similarity between the current layer’s query and the previous layers’ keys will be low, thus minimizing negative impacts.
-Given $m<n$ and the information $(\mathbf{Q}_{m},\mathbf{K}_{m},\mathbf{V}_{m})$ of $m$-th layer, the cross layer mechanism calculates the attention output $\mathbf{U}_{n}$ of $n$-th layer by the following attention formula:
-$\mathbf{U}_{n}$ $=\operatorname{Softmax}\Big{(}{\mathbf{Q}_{n}\operatorname{Concat}(\mathbf{K}_{n},\mathbf{K}_{m})^{T}}/{\sqrt{d}}\Big{)}\operatorname{Concat}(\mathbf{V}_{n},\mathbf{V}_{m}).$ (2)
-In practice, cross-layer attention enhances feature fusion by allowing information to flow between layers, capturing both intra-layer and inter-layer dependencies.
-However, this approach introduces additional computational overhead due to the concatenation of keys and values from multiple layers.
-For example, in scenarios described by Eqn.
-2, the overall computational complexity of the model nearly doubles compared with vanilla attention described in Eqn.
-1.
-3.2 Efficient Cross Layer Attention
-To solve this problem, we propose to replace the $\mathbf{K}_{m}$ with $\mathbf{K}_{n}$ in Eqn. 2, as shown in Eqn. 3.
-$\mathbf{U}_{n}$ $\approx\operatorname{Softmax}\Big{(}{\mathbf{Q}_{n}\operatorname{Concat}(\mathbf{K}_{n},\mathbf{K}_{n})^{T}}/{\sqrt{d}}\Big{)}\operatorname{Concat}(\mathbf{V}_{n},\mathbf{V}_{m})$ (3)
-$=\frac{1}{2}\operatorname{Softmax}\Big{(}{\mathbf{Q}_{n}\mathbf{K}_{n}^{T}}/{\sqrt{d}}\Big{)}({\mathbf{V}_{n}}+{\mathbf{V}_{m}}).$ (4)
-Utilizing the concept of block matrices, Eqn.
-3 can be further simplified into Eqn.
-4.
-This simplification converts the concatenation operation of the two value matrices into an addition operation.
-Compared to Eqn.
-1, this new method only brings a minimal increase in computational complexity while still leveraging the information from the $m$-th layer in the $n$-th layer.
-Furthermore, Eqn.
-4 can be generalized to incorporate cross-layer attention across all preceding $n-1$ layers as follows:
-$\mathbf{U}_{n}$ $\approx\frac{1}{n}\mathbf{A}_{n}\sum_{i=1}^{n}{\mathbf{V}_{i}}.$ (5)
-where $\mathbf{A}_{n}$ denotes the original attention matrix for layer $n$.
-From the perspective of information propagation, model described by Eqn.
-3 projects the historical values into the current layer’s embedding space using the current layer’s attention as a weight matrix.
-For example, a naive approach would be to perform identity mapping, as described by $\mathbf{U}_{n}$ $=\mathbf{A}_{n}\mathbf{V}_{n}+\frac{1}{n-1}\sum_{i=1}^{n-1}{\mathbf{V}_{i}}.$ (6) To evaluate the approximation effect of replacing the $\mathbf{K}_{m}$ with $\mathbf{K}_{n}$, we randomly select 1,000 pre-training data samples.
-For each layer of a trained baseline model, assuming cross-layer attention is required for each layer with respect to the previous one, we calculate the cosine similarity between the outputs from Eqn.
-2 and Eqn.
-4.
-We also calculate the cosine similarity between the outputs from Eqn.
-2 and Eqn.
-6 for comparison.
-Fig.
-4 shows that our proposed method provides a good approximation for cross-layer attention.
-[Uncaptioned image] Figure 4: Average token similarity between layer outputs.
-Lines show similarity of outputs using current attention (Eqn.
-4) or identity attention (Eqn.
-6) compared to the one using cross-layer attention (Eqn.
-2).
-3.3 Transformer with Residual Value
-Based on Eqn.
-5, we propose a variant of Transformer with residual value (ResFormer) which only chooses first layer as the target of cross layer attention since the first layer contains all basic information of each token.
-The analysis of entropy in Fig.
-1 (Right) supports this point, indicating that attention tends to be relatively dispersed across different tokens in the initial layers of the model.
-The attention mechanism of ResFormer can be formulated as
-$\mathbf{U}_{n}$ $=\frac{1}{2}\mathbf{A}_{n}({\mathbf{V}_{n}}+{\mathbf{V}_{1}}).$ (7)
-where $n\geq 2$ and standard attention is applied in the first layer.
-From the training perspective, it explicitly learns a residual mapping instead of directly learning the desired underlying mapping and that’s why we call it ResFormer.
-3.4 A unified View Of NeuTRENO and DenseFormer
-Using our framework, the NeuTRENO can be defined as
-$\mathbf{U}_{n}$ $=\Big{(}\mathbf{A}_{n}-\bm{\lambda}\mathbf{I}\Big{)}{\mathbf{V}_{n}}+\bm{\lambda}{\mathbf{V}_{1}}.$ (8)
-where $\mathbf{I}$ denotes the identity matrix and $\bm{\lambda}$ is a hyper-parameter.
-It can be found that the term of $\bm{\lambda}\mathbf{I}$ may have certain negative impact on the learning of original attention.
-If we ignore the attention output projection and the MLP layer, DenseFormer can also be modeled within our framework as
-$\mathbf{U}_{n}$ $=\sum_{i=1}^{n}{\bm{\alpha}_{i}\mathbf{A}_{i}\mathbf{V}_{i}}.$ (9)
-where $\{\bm{\alpha}_{i}\}_{i=1}^{n}$ is a set of hyper-parameters.
-DenseFormer uses attention matrix of previous layer as the weight matrix of projecting values but this is not aligned with the concept shown in Eqn.
-3.
-3.5 SVFormer: Single-layer Value For Half $KV$ Cache
-After ResFormer, a natural idea is whether we can remove the value vectors in each layer and have all layers share the value vectors from the first layer.
-We call this method SVFormer.
-Similar to ResFormer, SVFormer still adopts standard attention in the first layer and obtain the attention output $\mathbf{U_{n}}$ for $n$-th layer where $n\geq 2$ through
-$\mathbf{U}_{n}$ $=\mathbf{A}_{n}{\mathbf{V}_{1}}.$ (10)
-Compared to previous methods, SVFormer is the first method that decouple value vectors from attention.
-Its main advantage is that it only requires computing and storing the value vectors for the first layer, saving nearly half of the $KV$ cache during inference.
+For each layer, the hidden state $\mathbf{H}_{n-1}$ will be firstly projected into $\mathbf{Q}_{n},\mathbf{K}_{n},\mathbf{V}_{n}\in\mathbb{R}^{l\times d}$ through three linear projections $\mathbf{W^{\mathbf{Q}}_{n}},\mathbf{W^{\mathbf{K}}_{n}},\mathbf{W^{\mathbf{V}}_{n}}\in\mathbb{R}^{d\times d}$ respectively.
+After these projections, the attention operation ($\operatorname{Attn}$), output projection ($\mathbf{W}^{\mathbf{O}}_{n}\in\mathbb{R}^{d\times d}$), and Multi-Layer-Perceptron ($\operatorname{Mlp}$) are applied sequentially:
+$\mathbf{U}_{n}$ $=\operatorname{Attn}(\mathbf{Q}_{n},\mathbf{K}_{n},\mathbf{V}_{n}).$ (1)
+$\mathbf{H}_{n}$ $=\operatorname{Mlp}(\mathbf{U}_{n}\mathbf{W}^{\mathbf{O}}_{n}).$ (2)
+3.2 NeuTRENO and DenseFormer
+After Eqn. 1, NeuTRENO adds the difference between the first and current layer’s value:
+$\mathbf{U}_{n}$ $=\operatorname{Attn}(\mathbf{Q}_{n},\mathbf{K}_{n},\!\mathbf{V}_{n})+{\bm{\lambda}_{n}}(\mathbf{V}_{1}-\mathbf{V}_{n}).$ (3)
+After Eqn. 2, DenseFormer performs a weighted average between all previous hidden states:
+$\mathbf{H}_{n}$ $=\bm{\lambda}_{n,n}\operatorname{Mlp}(\mathbf{U}_{n}\mathbf{W}^{\mathbf{O}}_{n})+\sum_{i=1}^{n-1}\bm{\lambda}_{n,i}\mathbf{H}_{i}.$ (4)
+where $\mathbf{H}_{0}=\operatorname{Embedding}(\mathbf{X})$ for the input $\mathbf{X}$. $\bm{\lambda}_{n}$ in Eqn. 3 and $\!\{\bm{\lambda}_{n}\!\}$ in Eqn. 4 are new parameters.
+3.3 ResFormer
+In contrast, before Eqn.
+1, ResFormer introduces a skip connection from the first layer’s value $\mathbf{V}_{0}=\mathbf{H}_{0}\mathbf{W}^{\mathbf{V}}_{1}$ to current layer’s value $\mathbf{V}_{n}=\mathbf{H}_{n}\mathbf{W}^{\mathbf{V}}$:
+$\mathbf{V}_{n}$ $={\bm{\lambda}_{n,1}}\mathbf{V}_{1}+{\bm{\lambda}_{n,2}}\mathbf{H}_{n}\mathbf{W}^{\mathbf{V}}.$ (5)
+where ${\bm{\lambda}_{n,1}}$ and ${\bm{\lambda}_{n,2}}$ are flexible scalars.
+When all $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ are predetermined constants, it is termed Constant-ResFormer.
+If $\{{\bm{\lambda}_{n,1}}={\bm{\lambda}_{n,2}}\}_{n=1}^{N}$, where $N$ is the total number of layers, the model is called Identity-ResFormer.
+Another variant where some layers have $\bm{\lambda}_{1}=0$ is referred to as Sparse-ResFormer.
+Besides, if $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ are trainable parameters, the model is termed Learnable-ResFormer.
+Unless otherwise specified, $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ are initialized to 0.5 for Learnable-ResFormer and are predetermined as 0.5 for Identity-ResFormer.
+A more general form is the Dense-ResFormer, defined as $\mathbf{V}_{n}\!=\!\bm{\lambda}_{n,n}\mathbf{H}_{n}\mathbf{W}^{\mathbf{V}}+\sum_{i=1}^{n-1}\bm{\lambda}_{n,i}\mathbf{V}_{i}$ for $n\geq 2$, where $\{\bm{\lambda}_{n}\!\}$ are constants or trainable scalars.
+Unless noted, all $\bm{\lambda}$ are initialized to 1 here.
+3.4 SVFormer
+Shared Parts - values keys keys & values
+Valid Loss 2.739 2.743 2.753 2.776
+Table 1: Results of sharing different parts every 2 layers.
+Beyond ResFormer, SVFormer adopts standard attention in the first layer and obtain the attention output $\mathbf{U_{n}}$ for $n$-th layer where $n\geq 2$ through $\mathbf{U}_{n}=\mathbf{A}_{n}{\mathbf{V}_{1}}.$ Its main advantage is that it only requires computing and storing the value vectors for the first layer, saving nearly half of the $KV$ cache during inference.
 Similar methods like CLA reduce $KV$ cache by sharing both of the key and value vectors every two layers.
-However, the results in Fig.
-5 show that sharing values has less negative impact compared with sharing keys.
-Figure 5: Ablation study of sharing different parts of attention every two layers.
-4 Pretrain Experiments
+However, the results in Table  1 show that sharing values has less negative impact compared with sharing keys.
+4 Experiments
 4.1 Setting
-4.1.1 Training Details
+Training Details
 Following Brandon et al.
@@
 Specifically, the architecture includes pre-normalization, SwiGLU activations (Shazeer, 2020), rotary position embedding (Su et al., 2024), and no dropout.
-For slimpajama, we randomly sample nearly 20B tokens according to the original data distribution of seven domains during training and adopt tokenizer used for “RedPajama-INCITE-7B-Base”.
-The details of training data can be found in Table 2 in Appendix.
+For slimpajama, we randomly sample nearly 20B tokens based to the original data distribution of seven domains during training and adopt tokenizer used for “RedPajama-INCITE-7B-Base".
+See Table 12 in Appendix for data details.
 Unless otherwise noted, we train all models using AdamW optimizer with 0.1 weight decay, $\beta_{1}=0.9$, $\beta_{2}=0.95$ and the max grad norm 1.0.
@@
 We adopt linear learning rate warmup for the first 1,200 steps with the initial learning rate and the peak learning rate to be 1e-7 and 6e-4 respectively.
-The cosine decay schedule gradually decays to 10% of the peak learning rate by the end of training (Zhou et al., 2024, Wei et al., 2023).
+The cosine decay schedule gradually decays to 10% of the peak learning rate by the end of training (Zhou et al., 2024; Wei et al., 2023).
 The detailed hyperparameters for models of various sizes and different training sequence lengths in Appendix A.4.
@@
 We adopt deepspeed zero-2 optimizer and flash attention mechanism.
-4.1.2 Relative Training Loss Curve On SlimPajama
-We trained all models for only one epoch on SlimPajama subsets, and primarily use training loss to compare different models.
-Furthermore, we use the relative training loss curve for better visualizing the difference among different models from the perspective of loss landscape.
-Specifically, for each method, we will subtract the smoothed training curve of the vanilla Transformer, obtained under the same experimental settings, from the smoothed training curves of the method.
-The smoothing is done using a window size of 10 steps or 100 steps.
-4.1.3 Entropy For Analyzing Attention Concentration Effects
+Evaluation
+For model evaluation and comparison, we primarily utilized the average validation loss across seven domains, computed on the entire SlimPajama validation split.
+Additionally, we randomly selected a fixed set of 1,000 sample sequences for subsequent visualization analysis.
+4.2 Scaling Analysis of ResFormer
+Figure 2: (Left) Validation loss as model size scales from 82M to 468M parameters.
+(Right) Validation loss for the 468M parameter model evaluated every 2B tokens.
+ResFormer achieves approximately 13.3%-15.4% reduction in both model parameters and training data.
+We analyzed how ResFormer and Transformer scale at model size and data size.
+ResFormer and Transformer are trained on similar experiment setting.
+On the one hand, we trained model with 82M, 180M, 320M and 468M parameters on 20B training tokens and evaluated them on a separate validation set.
+As shown in Fig.2 (Left), ResFormer achieves equivalent validation loss to the Transformer while utilizing 13.3% fewer model parameters.
+On the other hand, we evaluated the 468M models every 2B tokens and ResFormer needs 15.4% fewer training tokens to achieve the same loss as Transformer.
+4.3 ResFormer vs. NeuTRENO, DenseFormer
+Model Initial Form Loss
+Baselines
+Vanilla Transformer - 2.739
+$\text{DenseFormer}^{*}$ $\mathbf{H}_{n}^{\prime}\!=\!{\color[rgb]{1,0,0}1\times}\mathbf{H}_{i}+\sum_{i=1}^{n-1}{\color[rgb]{1,0,0}0\times}\mathbf{H}_{i}$ 2.722
+NeuTRENO $\mathbf{U}_{n}^{\prime}={\color[rgb]{1,0,0}0.4}(\mathbf{V}_{1}-\mathbf{V}_{n})+\mathbf{U}_{n}$ 2.72
+ResFormer
+Identity-ResFormer $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}0.5}\mathbf{V}_{1}+{\color[rgb]{1,0,0}0.5}\mathbf{V}_{n}$ 2.712
+[…]
+[…]
+Constant-ResFormer $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}2}\mathbf{V}_{1}+{\color[rgb]{1,0,0}0.5}\mathbf{V}_{n}$ 2.7
+Sparse-ResFormer $\left\{\begin{array}[]{l}\mathbf{V}_{n}^{\prime}=\mathbf{V}_{n},1\leq n\leq 5\\ \mathbf{V}_{n}^{\prime}=\mathbf{V}_{1},6\leq n\leq 8\end{array}\right.$ 2.696
+Sparse-ResFormer $\left\{\begin{array}[]{l}\mathbf{V}_{n}^{\prime}=\mathbf{V}_{n},1\leq n\leq 5\\ \mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}5}\mathbf{V}_{1}+{\color[rgb]{1,0,0}0.5}\mathbf{V}_{n},\\ \qquad\ \quad\qquad 6\leq n\leq 8\end{array}\right.$ 2.687
+Table 2: Average valid loss for 8-layer, 82M-parameter models.
+“Initial form" shows deviations from vanilla transformer.
+Red numbers are the $\bm{\lambda}$ values from Eqn.
+3, Eqn.
+4, and Eqn.
+5.
+For models marked with “*" , $\bm{\lambda}$ is learnable, and the red numbers indicate the initial value; otherwise, red numbers are fixed constants.
+We compared the average validation loss of different models with 82M parameters trained on 20B tokens.
+In our experimental setup, NeuTRENO performed best with $\bm{\lambda}=0.4$, achieving comparable results to DenseFormer.
+All ResFormer variants, including the simplest Identity-ResFormer, demonstrated significant performance improvements.
+The $\bm{\lambda}$ values for Constant-ResFormer and Sparse-ResFormer were optimized through multiple experiments.
+While Sparse-ResFormer achieved the best overall performance, it is challenging to pre-determine the optimal layers for $\mathbf{V}_{1}$ connections and their corresponding $\bm{\lambda}_{1}$ values in more general scenarios.
+Although, Learnable-ResFormers are struggle to identify the optimal value residual pattern, they outperform Identity-ResFormer.
+Interestingly, the second to last row shows that Sparse-ResFormer achieved better performance despite having three fewer $\mathbf{W}^{\mathbf{V}}$.
+Figure 3: The impact of varying $\bm{\lambda}$ values on 82M 8-layer Constant-ResFormer and NeuTRENO.
+Both Constant-ResFormer and NeuTRENO rely on predetermined $\bm{\lambda}$ constants.
+Fig.
+3 shows the performance curves of these models against varying $\bm{\lambda}$.
+Results indicate that Constant-ResFormer significantly outperforms NeuTRENO and demonstrates greater robustness across a wider range of $\bm{\lambda}$ values, achieving optimal performance at $\bm{\lambda}=2$.
+4.4 Truly Better or Just Faster?
+Figure 4: (Left) Average gradient norms of model outputs with respect to parameter matrices across different layers in Transformer and ResFormer.
+(Right) Comparison of Transformer and ResFormer performance across various learning rates during training.
+To verify that ResFormer’s performance improvements are not solely due to accelerated training from its shortcuts, we examined model performance across different learning rates.
+We compared Identity ResFormer and Vanilla Transformer under five learning rate settings.
+As shown in Fig.
+4 (Right), both models achieved optimal results around a learning rate of 0.003, with Identity ResFormer significantly outperforming Vanilla Transformer across all rates.
+Analysis of the grad norm for the four parameter matrices ($\mathbf{W}^{\mathbf{Q}}$, $\mathbf{W}^{\mathbf{K}}$, $\mathbf{W}^{\mathbf{V}}$, $\mathbf{W}^{\mathbf{O}}$) in each layer’s attention module revealed that Identity ResFormer’s output had approximately twice the grad norm for $\mathbf{W}^{\mathbf{V}}_{1}$ and half for $\mathbf{W}^{\mathbf{O}}_{1}$ in the first layer compared to Vanilla Transformer.
+This indicates that a portion of the gradient originally propagated to $\mathbf{V}_{1}$ through $\mathbf{H}_{1}$ is now transmitted via the value residual directly for Identity ResFormer.
+In this way, we conducted the other two ablation experiments on Vanilla Transformer: doubling the learning rate for only the first layer, and doubling it exclusively for $\mathbf{W}^{\mathbf{V}}_{1}$ in the first layer.
+Neither modification yielded significant improvements.
+This further demonstrates that the performance improvements brought by ResFormer are unrelated to the changes in gradient magnitude.
+4.5 Ablation Study of Value Residual
+Where from, where to?
+Figure 5: (Left) Impact of value skip connections source from different layers on model performance, where all connections are identity connections and $\bm{\lambda}=1$ in Dense-ResFormer.
+(Right) Average validation loss of various Sparse-ResFormer configurations, which retain only single or multiple skip connections from $\mathbf{V}_{1}$.
+We analyzed which value skip-connections are necessary for the vanilla transformer.
+For an 8-layer transformer, we added various pre-defined value skip-connections (with constant $\bm{\lambda}$) and evaluated the resulting validation loss.
+As shown in Fig.
+5 (Left), we first examined the impact of skip-connections from different sources.
+Our findings indicate that only skip-connections originating from the first layer’s value ($\mathbf{V}_{1}$) yield significant performance improvements.
+Skip-connections from the second layer’s value ($\mathbf{V}_{2}$) offer no significant benefit to subsequent layers.
+Skip-connections from later layers, occurring only in the final few layers, even lead to performance degradation.
+Both of the two special cases in Fig.
+5 (Left) include $\mathbf{V}_{1}$ skip-connections.
+However, when these connections occur only between adjacent layers, the information in $\mathbf{V}_{1}$ fails to effectively reach the final layers.
+Conversely, dense value skip-connections dilute the impact of $\mathbf{V}_{1}$ with information from other sources.
+Furthermore, we investigated spare ResFormer, a variant of identity ResFormer where the value residual connection $\mathbf{V}_{n}^{\prime}={0.5}\mathbf{V}_{1}+{0.5}\mathbf{V}_{n}$ is applied selectively to specific layers.
+As shown in Fig.
+5 (Right), for an 8-layer model, when limited to a single layer, applying the residual connection to the 7th layer yields the most significant improvement.
+When applied to multiple layers, the greatest benefit is observed when incorporating layers 6 to 8.
+Extending the residual connection to earlier layers, such as the 5th, diminishes the overall effect.
+It suggests that the model’s final few layers benefit most from the first layer’s value information.
+Figure 6: (Left) Visualization of $\bm{\lambda}_{1}/\bm{\lambda}_{2}$ across different layers in the 82M and 468M Learnable-ResFormer.
+(Right) Heatmap visualization of learned $\bm{\lambda}$ across different layers in the 468M Dense-Learnable-ResFormer.
+We further trained 8-layer and 24-layer Learnable-ResFormers, as well as an 8-layer Learnable-Dense-ResFormer, and visualized the learned $\bm{\lambda}$ values.
+As shown in Fig.
+6, the later layers tend to require more value residual connections from $\mathbf{V}_{1}$, which aligns with the findings in Fig.
+5.
+Fortunately, the Learnable-ResFormer can, to some extent, identify similar sparse residual patterns to those of the best performing Sparse-ResFormer in Table 2.
+Notably, the Learnable-Dense-ResFormer learns value residual patterns that closely resemble those of the Learnable-ResFormer.
+Why needed beyond hidden residual?
+Residual Type Initial Form Valid Loss
+- - 2.7389
+value residual $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}0.5}\mathbf{V}_{1}+{\color[rgb]{1,0,0}0.5}\mathbf{V}_{n}$ 2.7049
+hidden residual $\mathbf{H}_{n}^{\prime}={\color[rgb]{1,0,0}0.5}\mathbf{H}_{0}+{\color[rgb]{1,0,0}0.5}\mathbf{H}_{n}$ 2.7812
+value residual $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}0\times}\mathbf{V}_{1}+{\color[rgb]{1,0,0}1}\mathbf{V}_{n}$ 2.7298
+hidden residual $\mathbf{H}_{n}^{\prime}={\color[rgb]{1,0,0}0\times}\mathbf{H}_{0}+{\color[rgb]{1,0,0}1}\mathbf{H}_{n}$ 2.7216
+Table 3: Comparison of additional value residual (to $\mathbf{V}_{1}$) and hidden residual (to $\mathbf{H}_{0}$) connections against the default hidden residual, under various $\mathbf{\lambda}$ initializations.
+Trainable $\mathbf{\lambda}$ parameters are highlighted in red.
+Our experiments revealed that $\mathbf{V}_{1}$ information provides additional benefits to later network layers, despite both $\mathbf{H}_{1}$ and $\mathbf{V}_{1}$ containing initial, unfused token information.
+$\mathbf{H}_{0}$ is propagated through default hidden residual connections, but it may be diluted by subsequent information, hindering its effective utilization in later layers.
+To test this hypothesis, we introduced an additional skip connection to $\mathbf{H}_{0}$ : $\mathbf{H}_{n}^{\prime}={\mathbf{\lambda}_{1}}\mathbf{H}_{1}+{\mathbf{\lambda}_{2}}\mathbf{H}_{n}$, where $\mathbf{\lambda}$ is learnable.
+We conducted experiments with two $\mathbf{\lambda}$ initialization settings and compared them to value residual.
+Results showed that when ${\mathbf{\lambda}_{1}}={\mathbf{\lambda}_{2}}$ initially, the extra hidden residual had adverse effects.
+However, initializing ${\mathbf{\lambda}_{1}}=0$ yielded some improvements, suggesting possible dilution of $\mathbf{H}_{0}$ information.
+Nevertheless, these gains were smaller than those from value residual connections, which consistently outperformed vanilla transformers across different initializations.
+Actually, the connection $\mathbf{H}_{0}$ : $\mathbf{H}_{n}^{\prime}={\mathbf{\lambda}_{1}}\mathbf{H}_{1}+{\mathbf{\lambda}_{2}}\mathbf{H}_{n}$, is similar to applying residuals to queries, keys, and values at the same time, may disrupt attention distributions and hinder higher-level semantic information fusing.
+The reduction performance brought by identity residuals of queries or keys shown in Table 6 can support it.
+Why $\mathbf{V}_{1}$ instead of $\mathbf{V}_{2}$?
+Hidden Residual Starts Place Value Residual Target Place Valid Loss
+$\mathbf{H_{0}}$ (Default) - 2.7389
+$\mathbf{H_{0}}$ (Default) $\mathbf{V_{1}}$ 2.7117
+$\mathbf{H_{0}}$ (Default) $\mathbf{V_{2}}$ 2.7375
+$\mathbf{H_{1}}$ 2.7802
+$\mathbf{H_{1}}$ $\mathbf{V_{2}}$ 2.784
+$\mathbf{H_{2}}$ 2.8196
+$\mathbf{H_{2}}$ $\mathbf{V_{2}}$ 2.7873
+$\mathbf{H_{2}}$ 2.8196
+$\mathbf{H_{2}}$ $\mathbf{V_{3}}$ 2.8333
+$\mathbf{H_{3}}$ 3.0568
+$\mathbf{H_{3}}$ $\mathbf{V_{3}}$ 2.8832
+Table 4: Comparison of performance across different value residual target and varying hidden residual start settings.
+“value residual target place" $\mathbf{V}_{i}$ indicates the earliest value accessible to subsequent layers, while “hidden residual starts place" denotes the earliest hidden state available, without prior residual connections.
+Figure 7: (Left) The relative training loss curve between different cross layer residual and vanilla hidden residual. (Right) Layer-to-layer hidden states similarity.
+In Fig.
+5 (Left), connections to $\mathbf{V}_{1}$ show significant improvement, while those to $\mathbf{V}_{2}$ yield minimal gains.
+This likely occurs because the original hidden residual propagates information from $\mathbf{H}_{1}$ to the network ($\mathbf{V}_{2}=\mathbf{H}_{1}\mathbf{W}^{\mathbf{V}}_{2}$).
+To verify, we adjusted residual connections, introducing them at different points.
+For example, starting from $\mathbf{H}_{1}$, we use $\mathbf{H}_{1}=Layer_{1}(\mathbf{H}_{0})$ instead of $\mathbf{H}_{1}=Layer_{1}(\mathbf{H}_{0})+\mathbf{H}_{0}$.
+Table 4 results show that when residual connections begin from $\mathbf{H}_{0}$ or $\mathbf{H}_{1}$, allowing $\mathbf{H}_{2}$ and subsequent layers access to $\mathbf{H}_{1}$, $\mathbf{V}_{2}+\mathbf{V}_{n}$ offers no improvement.
+However, starting from $\mathbf{H}_{2}$, skip connections from $\mathbf{V}_{2}$ provide substantial benefits.
+Regarding the disparity in information propagation between $\mathbf{V}_{2}$ (via $\mathbf{H}_{1}$) and $\mathbf{V}_{1}$ (via $\mathbf{H}_{0}$), we posit that after the first layer’s integration, $\mathbf{H}_{1}$ contains higher-level semantic information more similar to subsequent hidden states, see Fig.
+7 (Right).
+This may ensure that the attention distribution remains relatively undisturbed when connecting to $\mathbf{H}_{1}$.
+Besides, Fig.
+7 (Left) shows that Dense-ResFormer performs better than ResFormer when there is no cross layer hidden residual.
+Superior to other residual
+Residual Type Valid Loss
+- 2.739
+Query 2.742
+Key 2.746
+Attention 2.757
+Value 2.712
+Table 5: The impact of various residual types, where all residual connections adopt a form similar to $\mathbf{V}_{n}^{\prime}\!=\!{0.5}\mathbf{V}_{1}+{0.5}\mathbf{V}_{n}$.
+Residual Mapping Valid Loss
+- 2.739
+Identity Mapping 3.137
+Cross Layer Attention 2.729
+Current Attention 2.712
+Table 6: Comparison of different mapping matrices when adding $\mathbf{V}_{1}$ to $\mathbf{U}_{n}$, with “Current Attention" corresponding to Identity-ResFormer.
+For vanilla transformers, to better propagate information from the first layer, new residual connections can be introduced at various points in addition to the existing hidden residual: query states $\mathbf{Q}$, key states $\mathbf{K}$, value states $\mathbf{V}$, and post-softmax attention matrix $\mathbf{A}$.
+Results in Table 6 indicate that only the value residual connection improves performance.
+When connecting $\mathbf{V}_{1}$ and $\mathbf{V}_{n}$, three approaches free of extra parameters are possible: (1) the proposed residual connection, directly summing the two and then sharing an attention matrix; (2) cross layer attention (${\operatorname{Softmax}\Big{(}{\mathbf{Q}_{n}\operatorname{Concat}(\mathbf{K}_{n},\mathbf{K}_{1})^{T}}\Big{)}\operatorname{Concat}(\mathbf{V}_{n},\mathbf{V}_{1})}$), recomputing an attention matrix for $\mathbf{V}_{1}$ based on $\mathbf{K}_{1}$ and $\mathbf{Q}_{n}$; and (3) directly adding $\mathbf{V}_{1}$ to $\mathbf{U}_{n}$ in Eqn.
+1, equivalent to using an identity mapping as $\mathbf{V}_{1}$’s attention matrix in layer $N$.
+The second approach significantly increases computational cost.
+Results in Table 6 demonstrate that sharing the attention matrix yields the best performance.
+4.6 Post-Analysis of ResFormer
+Figure 8: (Left) Similarity between first layer values and other layers’ values. (Right) Token-to-token similarity across sequence for value states at different place.
+How value residual works?
+We performed post-analysis on trained ResFormer and vanilla Transformer models to understand value residual learning.
+Fig.
+8 (Left) shows cosine similarities between value states at different layers and the first layer, averaged across token positions.
+For ResFormer, we calculated this before and after applying value residual.
+Results show that in vanilla Transformers, the first layer’s value has low similarity with other layers.
+In contrast, ResFormer maintains high similarity between the first layer’s value and the post-residual values in subsequent layers due to value residual connections.
+Notably, in layers where ResFormer relies more heavily on the first layer’s value (see Fig.
+6), the pre-residual value exhibits lower similarity with the first layer’s value, indicating that $\mathbf{W}^{\mathbf{V}}$ in these layers is learning the value residual.
+For ResFormer, we also examined the average pairwise similarity between tokens’ values before and after the residual connection.
+The results Fig.
+8 (Right) reveal that with value residual connections, the learned values (before the value residual) from each layer become increasingly similar as the network deepens.
+We hypothesize that this is because, given the default hidden residual and value residual, each layer learns a $\Delta\mathbf{V}$, with the magnitude of necessary adjustments decreasing in later layers.
+This phenomenon is unique to ResFormer and not observed in vanilla Transformers.
+Figure 9: (Left) The change in test loss as model modules are progressively removed, starting from the back to front while keeping the first layer intact.
+(Right) The number of core features in each layer’s hidden state after PCA dimensionality reduction, where core features represent the minimum number of principal components required to explain $99\%$ of the variance.
+Representation and Module Analysis
+We analyzed the overall network changes, focusing on the hidden state representation capabilities and the contributions of different modules.
+Tyukin et al.
+(2024) suggests that removing $\operatorname{Attention}$ in Transformers has a significantly smaller impact than removing $\operatorname{Mlp}$.
+We progressively removed attention or MLP layers, starting from the last layer while retaining the first layer.
+Fig.
+9 (Left) demonstrates that for ResFormer, the impact of removing $\operatorname{Attention}$ is more comparable to that of removing $\operatorname{Mlp}$, in contrast to vanilla Transformers.
+This indicates that the $\operatorname{Attention}$ in ResFormer, with value residual, contribute more significantly to each layer’s hidden states than in vanilla Transformers.
+Furthermore, we performed PCA dimensionality reduction on the hidden states of each layer in both ResFormer and vanilla Transformer models.
+We determined the minimum number of principal components required to explain $99\%$ of the variance.
+Fig.
+9 reveals that ResFormer, starting from the second layer where value residual connections are introduced, consistently produces hidden states with a higher minimum number of principal components compared to vanilla Transformers.
+This suggests that ResFormer generates hidden states with higher information density.
+4.7 SVFormer vs. GQA,CLA
+Sequence Length Model Valid Loss
+2,048 - 2.739
+CLA2 2.776
+GQA2 2.748
+SVFormer 2.774
+64,000 - 2.753
+CLA2 2.793
+GQA2 2.773
+SVFormer 2.7485
+Sequence Length Model Valid Loss
+64,000 - 2.753
+GQA8 2.807
+CLA2   +GQA4 2.815
+SVFormer   +GQA4 2.741
+Table 7: Comparison of valid loss under varying degrees of $KV$ cache reduction.
+CLA2 denotes parameter sharing every two layers, while GQA2 indicates halving the key-value heads.
+Left: Model with nearly ${1}/{2}$ $KV$ cache.
+Right: Model with nearly ${1}/{8}$ $KV$ cache.
+Figure 10: Left: Relative training loss for SVFormer $vs.$ vanilla Transformer under different sequence lengths with a fixed batch size of 2M tokens.
+Right: Analysis of critical point, and we predict it for length 64,000 using linear regression with the last 1,000 data points.
+Architecture Learning Rate Warnup Steps Valid Loss
+Llama (82M) 1e-4 120 -0.033
+3e-4 120 +0.021
+6e-4 120 +0.035
+6e-4 1,200 +0.036
+GPT2 (78M) 6e-4 120 +0.029
+Table 8: Relative validation loss of SVFormer compared to vanilla Transformer under different hyper-parameter settings.
+Figure 11: Validation loss for SVFormer as model size scales from 82M to 468M.
+In the Table 7, at a training sequence length of 64,000, SVFormer demonstrates lower final loss compared to existing $KV$-efficient methods such as CLA and GQA.
+Moreover, it can be used concurrently with GQA to enhance $KV$ efficiency further.
+However, we observed that with a training sequence length of 2,048, SVFormer underperforms compared to GQA.
+The results indicate that sequence length significantly affects SVFormer’s performance.
+Thus, we conducted more comprehensive experiments on sequence length.
+Effects of sequence length
+Results in Fig.
+10 (Left) demonstrate that SVFormer will always be gradually surpassed by vanilla attention during training while its training speed is faster than vanilla Transformer at the early stage.
+However, as the training sequence length increases, the SVFormer model performs better.
+In this way, we focus on the critical point, defined as the number of training steps exceeded.
+Fig.
+10 (Right) illustrates that the relationship between the critical point and sequence length exhibits an exponential trend.
+We argue that it’s due to the challenge deep models face in fully optimizing the increasingly larger first-layer value matrix as the sequence length grows.
+Other factors
+Table 8 show SVFormer benefits more from smaller learning rates than from warmup.
+This aligns with performance correlating to total summed learning rate (Kaplan et al., 2020).
+Larger models, requiring smaller learning rates, suit SVFormer better.
+Results also indicates the SVFormer-Transformer difference is not architecture-sensitive.
+Compared with Transformer, SVFormer requires a 12.2% increase in parameters to achieve the same loss while reducing the $KV$-cache by nearly half in Fig.
+11.
+5 Conclusion
+In this paper, we demonstrate the inadequacy of existing hidden residual connections in propagating information from the initial token-level to deeper layers.
+To address this limitation, we propose ResFormer, which incorporates a residual connection between the value vectors of the current layer and those of the first layer prior to the attention operation.
+Furthermore, we introduce SVFormer, an extension of ResFormer, which achieves a nearly 50% reduction in the $KV$ cache.
+We conducted extensive experiments on language modeling tasks to evaluate the efficacy of these two Transformer variants across diverse scenarios.
+Limitations
+The proposed learnable ResFormer, still falls short of identifying the optimal $\bm{\lambda}$ setting through current training, instead converging on a relative optimum.
+This limitation suggests that further refinement of initialization strategies and learning algorithms may be necessary.
+Due to computational constraints, we were unable to conduct experimental validation on larger-scale models at this time.
+Ethics Statement
+On the one hand, the data employed in this paper is sourced from publicly available datasets provided by the company, which have undergone a certain level of filtering.
+On the other hand, the models trained in our study are solely utilized for experimental analysis and will not be publicly deployed.
+Appendix A Appendix
+A.1 Attention pattern analysis
+Figure 12: (Left) Average entropy of token importance and the average hidden-state similarity for a randomly initialized 468M model.
+(Right) Average entropy of token importance across layers in Llama (8B) (Dubey et al., 2024) and Mistral (7B) (Jiang et al., 2023).
+(a) Attention Entropy.
+(b) Token importance.
+(c) Norms of value states.
+(d) Norms of hidden states.
+Figure 13: The token importance (Xiao et al., 2024), value-state norms (Guo et al., 2024b), and hidden-state norms (Sun et al., 2024a) of the first token across layers of 468M models.
+“Attention Entropy" refers to the entropy of token importance across each sequence.
+Attention concentration
 Given the attention matrix $\mathbf{A}\in\mathbb{R}^{l\times l}$ at one layer, we use entropy $\bm{e}$ to represent its concentration effect.
 To obtain entropy $\bm{E}$, calculate the importance vector $\mathbf{a}=\frac{1}{l}\sum_{j=1}^{l}A_{ij}$ firstly where $\mathbf{A}$ is a lower triangular matrix.
-The entropy can be formulated as follows:
-$\bm{e}$ $=-\sum_{i=1}^{l}\bm{a}_{i}^{\prime}\log{\bm{a}^{\prime}_{i}}.$ (11)
-where $a_{i}^{{\prime}_{i}}={a_{i}}/\Big{(}{\sum_{i=1}^{l}{a_{i}}}\Big{)}$ for $i=1,2,\dots,l$ and the higher the entropy $\bm{e}$, the greater the degree of clustering in $\bm{a}$, i.e., attention matrix $\mathbf{A}$ is more likely to focus on several specific tokens.
-4.1.4 Spectral Decomposition For Analyzing Representations
-Spectral Decomposition is a classical method to analyze the representations of models.
-Zhu et al.
-(2021) suggests that the eigenvectors with larger eigenvalues are more transferable.
-Here we use spectral decomposition to analyze the feature space of value $\bm{v}$ of one layer as following:
-$\frac{1}{l}\sum_{i=1}^{l}\bm{v}_{i}\bm{v}_{i}^{T}$ $=\sum_{j=1}^{d}\bm{w}_{j}\bm{\lambda}_{j}\bm{w}_{j}^{T}.\vspace{-3mm}$ (12)
-where $\bm{w}_{j}$ is the $j$-th eigenvector with eigenvalue $\bm{\lambda}_{j}$ for $j=1,2,\dots,d$ and $d$ is the dimensionality of the value’s feature space.
-Figure 6: (Left) Validation loss as model size scales from 82M to 468M parameters.
-(Right) Validation loss for the 468M parameter model evaluated every 2B tokens.
-ResFormer achieves approximately 10.4%-13.6% reduction in both model parameters and training data.
-Figure 7: (Left) The relative training curve between a 82M ResFormer and Transformer across different training sequence lengths.
-(Right) Relative training loss of various Transformer variants compared to the vanilla Transformer model, with model size fixed at 82M parameters.
-4.2 ResFormer vs. vanilla Transformer
-We trained ResFormer and vanilla Transformer with different model size on data with different sequence lengths.
-In Fig.
-7 (Left), ResFormer consistently outperforms vanilla Transformer throughout training across different training sequence lengths.
-Additionally, the results in Fig.
-7 (Right) illustrate that ResFormer outperforms DenseFormer and NeuTRENO.
-Furthermore, integrating ResFormer with NeuTRENO leads to additional performance improvements.
-We also analyzed how ResFormer and Transformer scale at model size and data size.
-ResFormer and Transformer are trained on similar experiment setting.
-On the one hand, we trained model with 82M, 180M, 320M and 468M parameters on 20B training tokens and evaluated them on a separate validation set.
-As shown in Fig.7 (Left), ResFormer achieves equivalent validation loss to the Transformer while utilizing 10.4% fewer model parameters.
-On the other hand, we evaluated the 468M models every 2B tokens and ResFormer needs 13.6% fewer training tokens to achieve the same validation loss as Transformer.
-The validation loss for these models is available in Appendix A.6.
-We further test the variant of ResFormer defined as $\mathbf{U}_{n}=\mathbf{A}_{n}({\mathbf{V}_{n}}+{\bm{\lambda}}{\mathbf{V}_{1}})$.
-As shown in Fig.8, ResFormer can accommodate a wide range of $\bm{\lambda}$ values and the performance improves as $\bm{\lambda}$ increases, achieving the best results at ${\bm{\lambda}}=2$.
-Regardless of the value of $\bm{\lambda}$, ResFormer consistently outperforms Transformers.
-It suggests that the success of ResFormer lies in the use of $\mathbf{V}_{1}$ and the mapping by $\mathbf{A}_{n}$.
-The ablation study of different hyperparameters $\bm{\lambda}$ for NeuTRENO, as defined in Equation 8, can be found in the Appendix A.3.
-Figure 8: Ablation study of different $\bm{\lambda}$ for ResFormer.
-Figure 9: Ablation study of adding residual connection to queries or keys.
-Figure 10: Ablation study of adding residual connection using different mapping matrix.
-Figure 11: Ablation studies on which historical layer’s value to include in residual connections.
-4.3 Ablation study of residual connection
-In Eqn.
-4, we employ residual connections for the values.
-We compare this approach with models that add residual connections to queries or keys.
-The results, shown in Fig.
-11, indicate that only residual connections for values yield positive effects.
-One possible explanation is that attention mechanisms are sensitive to perturbations, and modifying queries or keys significantly impacts it.
-Moreover, we compare with the models based on Eqn.
-2 and Eqn.
-6.
-The results in Fig.
-11 align with Fig.
-4, showing that identity mapping causes significant perturbations, leading to poor performance.
-Interestingly, ResFormer achieves an even lower final loss than ResFormer.
-It suggests that ResFormer’s impact on the attention optimization is better by mitigating value-state drains.
-When determining the mapping method and target value, it is crucial to consider which historical layers’ values should be included in the residual connection.
+The entropy can be formulated as follows: $\bm{e}=-\sum_{i=1}^{l}\bm{a}_{i}^{\prime}\log{\bm{a}^{\prime}_{i}}$, where $a_{i}^{{\prime}}={a_{i}}/\Big{(}{\sum_{i=1}^{l}{a_{i}}}\Big{)}$ for $i=1,2,\dots,l$ and the higher the entropy $\bm{e}$, the greater the degree of clustering in $\bm{a}$, i.e., attention matrix $\mathbf{A}$ is more likely to focus on specific tokens.
+The phenomenon of attention concentration is inherent to model architecture and emerges during training.
 Fig.
-11 shows that each Transformer layer should add a shortcut to the first layer’s value rather than to the nearest preceding layer or all previous layers, highlighting the first-layer value’s critical importance.
-A potential explanation is that incorporating values from other layers may dilute the impact of the first-layer value.
-4.4 Downstream Evaluations
-We compare the different models on several classical reasoning tasks following (Zhang et al., 2024) in a zero-shot way.
+12 shows that randomly initialized models exhibit over-smoothing but not attention concentration and popular trained models exhibit obvious attention concentration problem.
+Trained ViT models often focus on low-informative background areas (Darcet et al., 2024), while language models concentrate on low-semantic tokens (Sun et al., 2024a), particularly the start token (attention sink (Xiao et al., 2024)).
+While previous studies analyzed single-layer attention patterns, our research reveals a “concentration - dispersion - concentration" pattern in deep models, as shown in Fig.
+12 (Right), suggesting potential loss of information during concentrated phases.
+Resformer alleviates attention concentration
+Transformer
+NeuTRENO
+ResFormer
+SVFormer
+(e) Token importance.
+(f) Value-state norms.
+(g) Hidden-state drains.
+Figure 14: Visualization of token importance, value state norms, and hidden state norms across different token positions and layers in 468M models.
+(a) Transformer.
+(b) NeuTRENO.
+(c) ResFormer.
+(d) SVFormer.
+Figure 15: The distribution of token importance for different models at different layers.
+Fig.
+13(a) illustrates that the clustering effect of attention increases significantly with the number of layers for the vanilla Transformer, whereas the clustering effect is relatively less pronounced for the ResFormer.
+We further visualize the attention weights, value-state norms $\|\bm{v}\|_{2}$, and hidden-state norms $\|\bm{h}\|_{2}$ of tokens at different layers and positions.
+Given that attention clustering often occurs on the first token, we primarily show its results in Fig.
+13.
+The results indicate that using ResFormer significantly mitigates attention sinks (Xiao et al., 2024), value-state drains (Guo et al., 2024b) and residual-state peaks (Sun et al., 2024a).
+Guo et al.
+(2024a) attributes these phenomena to the mutual reinforcement mechanism of model between value-state drains and attention sinks.
+We suggest that the value shortcut disrupts this mechanism by alleviating value-state drains due to the absence of value-state drains in the first layer.
+Specifically, for tokens lacking semantic information like start tokens, a large value state magnitude can adversely affect the prediction of subsequent tokens if they are overly attended to since $\mathbf{U}_{n}=\mathbf{A}_{n}\mathbf{V}_{n}$ in Eqn.1.
+However, when there is no value-state drains, models will reduce attention clustering to these tokens to minimize loss.
+Fig.
+14 (First column) demonstrates that the start token easily attracts massive attention despite lacking semantic information for Transformer and NeuTRENO.
+And Fig.15 further illustrates the distribution of token importance, where TOP-$i$ represents the $i$-th largest token importance within a sequence.
+Compared to Transformer and NeuTRENO, ResFormer and SVFormer exhibit a more uniform distribution of token importance.
+Any negative effect?
+$\lambda$ value Layers Valid loss
+- - 2.739
+1 All layers 2.7037
+2,3,4 2.724
+5,6,7 2.698
+2 All layers 2.7
+2,3,4 2.728
+5,6,7 2.688
+5 All layers 2.704
+2,3,4 2.732
+5,6,7 2.682
+Table 9: Performance of Sparse Resformer with different $\lambda$ values and different layer configurations.
+ResFormer Layers Norm re-scale Valid loss
+All Layers No 2.712
+Yes 2.701
+2,3,4 No 2.734
+Yes 2.727
+6,7,8 No 2.702
+Yes 2.701
+Table 10: Ablation study of post-value residual re-scaling.
+Mitigating attention concentration may enhance interpretability but potentially affect transformer sparsity.
+As attention sinks typically emerge early, applying value residual to later layers should have less impact intuitively.
+We compared two sparse Resformer variants on an 8-layer model, applying value residual to layers 2-4 versus 6-8.
+The results in Table 9 demonstrate that while incorporating value residual generally improves performance compared to vanilla Transformers, increasing $\lambda$ (the proportion of $\mathbf{V}_{1}$) in the value residual led to decreased performance for shallower networks.
+Conversely, deeper networks showed improved results with higher $\lambda$ values.
+Notably, the Learnable-ResFormer learned to apply value residual primarily to later layers, minimizing the impact on network sparsity.
+Moreover, we implemented post-value residual re-scaling ($\mathbf{V}_{n}^{\prime}=\frac{\|\mathbf{V}_{n}\|}{\|0.5\mathbf{V}_{1}+0.5\mathbf{V}_{n}\|}({0.5}\mathbf{V}_{1}+{0.5}\mathbf{V}_{n})$) to mitigate its impact on later layers.
+This approach benefited shallow-sparse Resformers but had minimal effect on deep-sparse variants.
+This further suggests that the value residual patterns learned by the Learnable ResFormer do not introduce significant negative effects in this context.
+A.2 Downstream Evaluations
+We compare the ResFormer and vanilla Transformer on several classical reasoning tasks following (Zhang et al., 2024) in a zero-shot way.
 The tasks include Hellaswag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2019), ARC-Easy and ARC-Challenge (Clark et al., 2018) and PIQA (Bisk et al., 2020).
@@
 Table 11: Zero-shot accuracy on commonsense reasoning tasks.
-Figure 12: Left: Distribution of eigenvalues for the value vectors in the first layer of ResFormer and Transformer. Right: Maximum eigenvalue for each layer of ResFormer and Transformer.
-4.5 Visualization of ResFormer
-To figure out why ResFormer can achieve better performance on language modeling tasks than vanilla Transformer, we conduct visualization based on the eigenvalue decomposition discussed in Section 4.1.4.
-After sorting the eigenvalues in descending order, we compute the average eigenvalue for each layer across 1,000 randomly sampled pre-train data examples.
-The results in Fig.
-12 indicate that the value states generated by most layers of the ResFormer exhibit stronger representational capacity compared to those of the vanilla Transformer.
-We also analyze the attention concentration effects mentioned in Section 4.1.3 using the same batch of test data.
-Fig.
-1 (Right) illustrates that the clustering effect of attention increases significantly with the number of layers for the vanilla Transformer, whereas the clustering effect is relatively less pronounced for the ResFormer.
-We further visualize the attention weights, value-state norms $\|\bm{v}\|_{2}$, and hidden-state norms $\|\bm{h}\|_{2}$ of tokens at different layers and positions, with detailed results in Appendix A.2.
-Given that attention clustering often occurs on the first token, we primarily show its results in Fig.
-2.
-The results indicate that using ResFormer significantly mitigates attention sinks (Xiao et al., 2024), value-state drains (Guo et al., 2024b) and residual-state peaks (Sun et al., 2024).
-Guo et al.
-(2024a) attributes these phenomena to the mutual reinforcement mechanism of model and we suggest that the value shortcut disrupts this mechanism by alleviating value-state drains.
-Specifically, for tokens lacking semantic information like start tokens, a large value state magnitude can adversely affect the prediction of subsequent tokens if they are overly attended to.
-When there is no value-state drains, models will reduce attention clustering to these tokens to minimize loss.
-4.6 SVFormer vs. GQA
-Figure 13: The relative training loss for SVFormer and other $KV$ efficient model compared with vanilla attention.
-The numbers in parentheses represent the training sequence length.
-Left: Model with nearly ${1}/{2}$ $KV$ cache.
-Right: Model with nearly ${1}/{8}$ $KV$ cache.
-Figure 14: Left: The relative training loss for SVFormer under different sequence lengths with a fixed batch size of 2M tokens.
-Right: Analysis of critical point, and we predict it for length 64,000 using linear regression with the last 1,000 data points.
-In the Fig.
-14, at a training sequence length of 64,000, SVFormer demonstrates lower final loss compared to existing $KV$-efficient methods such as CLA and GQA.
-Moreover, it can be used concurrently with GQA to enhance $KV$ efficiency further.
-However, we observed that with a training sequence length of 2,048, SVFormer underperforms compared to GQA.
-The results indicate that sequence length significantly affects SVFormer’s performance.
-Thus, we conducted more comprehensive experiments on sequence length.
-Results in Fig.
-14 (Left) demonstrate that SVFormer will always be gradually surpassed by vanilla attention during training while its training speed is faster than vanilla Transformer at the early stage.
-However, as the training sequence length increases, the SVFormer model performs better.
-In this way, we focus on the critical point, defined as the number of training steps exceeded.
-Fig.
-14 (Right) illustrates that the relationship between the critical point and sequence length exhibits an exponential trend.
-We argue that it’s due to the challenge deep models face in fully optimizing the increasingly larger first-layer value matrix as the training sequence length grows.
-4.7 Other Factors Influencing SVFormer
-Intuitively, the training effectiveness of SVFormer is influenced by factors such as the maximum learning rate, warmup steps, model size, and other factors beyond just the training sequence length.
-We conducted experiments to explore these relationships.
-Based on the results shown in Fig.
-15(a) and Fig.
-15(b), a smaller learning rate benefits SVFormer more, with warmup’s impact being comparatively small.
-This could be attributed to the model’s outcomes being closely tied to the total summed learning rate, which has weak connection with warmup steps (Kaplan et al., 2020).
-Moreover, larger models often require smaller learning rates to ensure training stability, making them more suitable for using SVFormer.
-Llama-like models and GPT2-like models exhibit similar critical points and final losses (see Fig.
-15(c)).
-This suggests that the difference between SVFormer and the vanilla Transformer is not sensitive to architecture.
-Compared with Transformer, SVFormer requires a 12.2% increase in parameters to achieve the same validation loss while reducing the $KV$-cache by nearly half.
-(a) Learning Rate.
-(b) Warmup Steps.
-(c) Architecture.
-(d) Model Size.
-Figure 15: The relative training loss for SVFormer under different hyper-parameter setting and the validation loss as model size scales from 82M to 468M parameters.
-4.8 Ablation study of SVFormer
-Figure 16: Ablation study of sharing first layer’s query(key) across all layers.
-Figure 17: Ablation study on sharing values from different numbers of layers.
-To better understand SVFormer, we conduct several ablation experiments.
-We first observe the effects of sharing the first layer’s queries or keys across all layers in Fig.
-17, finding that this significantly impacts model performance, similar to the results in Fig.
-5.
-Additionally, sharing the first layer’s values in a multi-layer network may reduce the network’s “effective depth.” By updating the shared values using intermediate layers as “anchors,” we find that increasing the number of “anchors” improves performance, as shown in Fig.
-17.
-5 Conclusion
-In this paper, we propose the concept of attention concentration, a problem that arises from stacking multiple attention layers.
-From the perspective of cross-layer attention, we derive ResFormer, which adds a residual connection between the value vectors of the current layer and those of the first layer before the attention operation to alleviate attention concentration.
-Additionally, we introduce SVFormer, based on ResFormer, which reduces the $KV$ cache by nearly half.
-We conducted comprehensive experiments on the language modeling task to validate the advantages of these two Transformer variants in different scenarios.
-Ethics Statement
-On the one hand, the data employed in this paper is sourced from publicly available datasets provided by the company, which have undergone a certain level of filtering.
-On the other hand, the models trained in our study are solely utilized for experimental analysis and will not be publicly deployed.
-Reproducibility Statement
-We have detailed the complete experiments setup such as batch size, optimizer, learning rates in Section 4.1.1.
-Besides, we will release source codes once our paper is made public.
-These resources should be sufficient to reproduce results of the paper.
-Appendix A Appendix
-A.1 Token Similarity Analysis
-Attention concentration tends to make embeddings of different tokens more similar, resulting in over-smoothing.
-The extent of over-smoothing can be assessed by calculating the average token similarity $\bm{s}$ of the hidden states using the following formula:
-$\bm{s}$ $=\frac{1}{l(l-1)}\sum\nolimits_{i=1}^{l}\sum\nolimits_{j=1,j\neq i}^{l}\operatorname{Sim}\Big{(}\bm{h}_{i},\bm{h}_{j}\Big{)}.$ (13)
-where $\{\bm{h}_{i}\}_{i=1}^{l}$ is the hidden state of the $i$-th token and $\operatorname{Sim}(\cdot)$ denotes the operation of cosine similarity.
-The results in Fig.
-18 are align with the results in Fig.
-1.
-In the case of Llama and Mistral, the average token similarity demonstrates an “M”-shaped pattern with increasing network depth, while entropy follows a “W”-shaped pattern at corresponding positions.
-These trends indicate a strong correlation between attention concentration and over-smoothing.
-Figure 18: The average token similarity of hidden states across layers in Llama and Mistral.
-A.2 Attention Concentration Visualization
-Transformer
-NeuTRENO
-ResFormer
-SVFormer
-(a) Token importance.
-(b) Value-state norms.
-(c) Hidden-state drains.
-Figure 19: Visualization of token importance, value state norms, and hidden state norms across different token positions and layers in 468M models.
-(a) Entropy.
-(b) Token importance.
-(c) Value-state norms.
-(d) Hidden-state norms.
-Figure 20: Token analysis in 82M models. (a) Importance entropy of sequences; (b) Token importance, (c) value-state norms, and (d) hidden-state norms of the first token across layers.
-(a) Transformer.
-(b) NeuTRENO.
-(c) ResFormer.
-(d) SVFormer.
-Figure 21: The distribution of token importance for different models at different layers.
-We visualize the token importance, norms of value states and norms of hidden states for tokens at different position across layers.
-The results are averaged from 1,000 different sequences so that only the start token is the same and special across all sequences.
-Fig.
-19 (First column) demonstrates that the start token easily attracts massive attention despite lacking semantic information for Transformer and NeuTRENO.
-For ResFormer, the importance of the start token is less than 10 times that of tokens at other positions, indicating that tokens carrying semantic information receive more attention.
-Moreover, both Transformer and NeuTRENO exhibit significant value-state drains (Guo et al., 2024b) and residual-state peaks (Guo et al., 2024a, Sun et al., 2024) on the start token at certain layers.
-In contrast, for ResFormer, the value state norm of the start token exceeds half the magnitude of other tokens, while the peak hidden state norm is less than triple the average.
-Fig.21 further illustrates the distribution of token importance, where TOP-$i$ represents the $i$-th largest token importance within a sequence.
-Compared to Transformer and NeuTRENO, ResFormer and SVFormer exhibit a more uniform distribution of token importance.
-Similar to Fig.2, we conducted experiments on 82M models, with results shown in Fig.20.
-We also illustrate the attention pattern of cross-layer attention introduced in Eqn.
-2.
-The results demonstrate that while cross-layer attention successfully mitigates the problem of attention concentration, it still exhibits value-state drains.
-A.3 Ablation study of NeuTRENO
-NeuTRENO is sensitive to the choice of hyperparameter $\bm{\lambda}$ which is task-dependent.
-In the appendix of Nguyen et al.
-(2023), it is reported that $\lambda$ is set to 0.6 for image classification and segmentation tasks, and 0.4 for language modeling tasks.
-Fig.
-22 indicates that ${\bm{\lambda}}=0.4$ achieves the best results in our training dataset so that we choose ${\bm{\lambda}}=0.4$ for comparison.
-Besides, we empirically choose ${\bm{\lambda}}=0.2$ for NeuTRENO when combined with ResFormer.
-Figure 22: Ablation study of different $\lambda$ for NeuTRENO.
-A.4 Pre-train Dataset
+A.3 Pre-train Dataset
 Data source proportions Tokens
@@
 In this way, we sampled 20B tokens data from original 627B data based on the original data proportions shown in the Table 12.
-A.5 Training Details
+A.4 Training Details
 Max Sequence Length 512 2,048 8,192 32,000 64,000
@@
 Table 13: Training details for training dataset with different sequence length.
-Section 4.1.1 introduces the main experimental hyperparameters used in the paper.
+Section 4.1 introduces the main experimental hyperparameters used in the paper.
 This section further details the training parameters for various model sizes and training sequence lengths.
@@
 Table 14: Training details for models with different size.
-A.6 Validation Loss On Slimpajama
-Section 4.1.2 introduces to use relative training loss as a main evaluation matrix. Table 5 reports the validation loss for differnt model on the whole validation split of slimpajama.
-Model
-Common
-Crawl
-C4 Github
-Stack
-Exchange
-Wikipedia Book Arxiv Avg.
-Transformer (82M) 3.3595 3.5388 1.4247 2.3872 2.9047 3.3797 2.1779 2.7389
-Transformer (180M) 3.0961 3.2834 1.2451 2.1651 2.5897 3.1309 2.0001 2.5015
-Transformer (468M) 2.8514 3.0430 1.0908 1.9628 2.2821 2.8979 1.8362 2.2806
-ResFormer (82M) 3.3362 3.5191 1.3941 2.3592 2.8646 3.3572 2.1518 2.7117
-ResFormer (180M) 3.0631 3.2504 1.2200 2.1350 2.5435 3.0994 1.9732 2.4692
-ResFormer (468M) 2.8214 3.0115 1.0730 1.9388 2.2477 2.8696 1.8142 2.2537
-Table 5: Validation loss on slimpajama.

### v5

@@
 This paper introduces ResFormer, a novel architecture that enhances information flow by incorporating value residual connections in addition to hidden state residuals.
-And a variant is the SVFormer, where all layers share the first layer’s value embedding.
-Comprehensive empirical evidence demonstrates ResFormer achieves equivalent validation loss with 13.3% fewer model parameters and 15.4% less training data compared to Transformer, while maintaining similar memory usage and computational cost.
+And a variant is SVFormer, where all layers share the first layer’s value embedding.
+Comprehensive empirical evidence demonstrates ResFormer achieves equivalent validation loss with 16.11% fewer model parameters and 20.3% less training data compared to Transformer, while maintaining similar memory usage and computational cost.
 Besides, SVFormer reduces KV cache size by nearly half with only a small performance penalty and can be integrated with other KV-efficient methods, yielding further reductions in KV cache, with performance influenced by sequence length and cumulative learning rate.
 Value Residual Learning
-Zhanchao Zhou1,2   Tianyi Wu3†thanks: Equal Contribution; †Work done during internship at Westlake University; $\diamond$ Corresponding author.
-Zhiyun Jiang4†∗   Fares Obeid5   Zhenzhong Lan2⋄ 1Zhejiang University      2Westlake University 3University of Electronic Science and Technology of China 4China University of Mining and Technology     5Imperial College London
+Zhanchao Zhou1,2,3   Tianyi Wu4†thanks: Equal Contribution; †Work done during internship at Westlake University; $\diamond$ Corresponding author.
+Zhiyun Jiang5†∗   Fares Obeid6   Zhenzhong Lan2⋄ 1Zhejiang University      2Westlake University      3Ant Group 4University of Electronic Science and Technology of China 5China University of Mining and Technology     6Imperial College London
 1 Introduction
@@
 In a standard decoder-only transformer, initial token embeddings contain localized information, which rapidly evolves into abstract semantic features through early attention layers (Sun et al., 2024b; Clark et al., 2019).
-As Transformer deepens, a critical question arises: How effectively is the initial information propagated to deeper layers?
+As Transformers deepen, a critical question arises: How effectively is the initial information propagated to deeper layers?
 One common answer is that residual connections of hidden states ensure access to initial information throughout the network.
@@
 This indicates that in deeper layers, sequence-level features become dominant, while token-level features are diluted.
-DenseFormer (Pagliardini et al., 2024) applied the idea of learnable dense connections from DenseNet (Huang et al., 2016) to transformers, and the learned connection coefficients shows that deeper layers indeed require larger attention to initial embeddings.
+DenseFormer (Pagliardini et al., 2024) applied the idea of learnable dense connections from DenseNet (Huang et al., 2016) to Transformer, and the learned connection coefficients shows that deeper layers indeed require larger attention to initial embeddings.
 Given the low similarity between initial token embeddings and deeper hidden states (Sun et al., 2024b), their directly summation may significantly impact the modeling of attention distribution for abstract semantic information in later layers.
@@
 Existing $KV$-efficient methods often process keys and values simultaneously.
-Building on ResFormer, we decouple the value from the attention operation and propose a new kind of Transformer (SVFormer) where all layers share a single value states.
+Building on ResFormer, we decouple the value from the attention operation and propose a new kind of Transformer (SVFormer) where all layers share a single value state.
 We experiment on a 20B SlimPajama sub-sampled dataset, using settings similar to popular large language models (Wei et al., 2023; Dubey et al., 2024; Kaplan et al., 2020).
@@
 Results show that ResFormer outperforms the vanilla Transformer, DenseFormer, and NeuTRENO.
-ResFormer achieves equivalent validation loss with 13.3% fewer model parameters and 15.4% less training data compared to Transformer, while maintaining similar memory usage and computational cost.
+ResFormer achieves equivalent validation loss with 16.11% fewer model parameters and 20.3% less training data compared to Transformer, while maintaining similar memory usage and computational cost.
 Besides, SVFormer, while reducing the $KV$-cache by nearly half, requires a 12.2% increase in parameters to achieve the same validation loss as Transformer.
@@
 3 Method
-3.1 Notations
-Let $\mathbf{H}_{n}\in\mathbb{R}^{l\times d}$ be the input hidden state of the $n$-th layer, where $l$ denotes the sequence length and $d$ is the dimension size.
+3.1 Preliminary
+Notations
+Let $\mathbf{H}_{n}\in\mathbb{R}^{l\times d}$ be the output hidden state of the $n$-th layer, where $l$ denotes the sequence length and $d$ is the dimension size.
 For each layer, the hidden state $\mathbf{H}_{n-1}$ will be firstly projected into $\mathbf{Q}_{n},\mathbf{K}_{n},\mathbf{V}_{n}\in\mathbb{R}^{l\times d}$ through three linear projections $\mathbf{W^{\mathbf{Q}}_{n}},\mathbf{W^{\mathbf{K}}_{n}},\mathbf{W^{\mathbf{V}}_{n}}\in\mathbb{R}^{d\times d}$ respectively.
@@
 $\mathbf{H}_{n}$ $=\operatorname{Mlp}(\mathbf{U}_{n}\mathbf{W}^{\mathbf{O}}_{n}).$ (2)
-3.2 NeuTRENO and DenseFormer
+NeuTRENO and DenseFormer
 After Eqn. 1, NeuTRENO adds the difference between the first and current layer’s value:
@@
 After Eqn. 2, DenseFormer performs a weighted average between all previous hidden states:
-$\mathbf{H}_{n}$ $=\bm{\lambda}_{n,n}\operatorname{Mlp}(\mathbf{U}_{n}\mathbf{W}^{\mathbf{O}}_{n})+\sum_{i=1}^{n-1}\bm{\lambda}_{n,i}\mathbf{H}_{i}.$ (4)
-where $\mathbf{H}_{0}=\operatorname{Embedding}(\mathbf{X})$ for the input $\mathbf{X}$. $\bm{\lambda}_{n}$ in Eqn. 3 and $\!\{\bm{\lambda}_{n}\!\}$ in Eqn. 4 are new parameters.
+$\mathbf{H}_{n}$ $=\bm{\lambda}_{n,n}\operatorname{Mlp}(\mathbf{U}_{n}\mathbf{W}^{\mathbf{O}}_{n})+\sum_{i=0}^{n-1}\bm{\lambda}_{n,i}\mathbf{H}_{i}.$ (4)
+where $\mathbf{H}_{0}=\operatorname{Embedding}(\mathbf{X})$ for the input $\mathbf{X}$.
+$\bm{\lambda}_{n}$ in Eqn.
+3 and $\!\{\bm{\lambda}_{n,i}\!\}_{i=0}^{n-1}$ in Eqn.
+4 are new parameters.
+Unless noted, $\bm{\lambda}_{n}$ is set to 0.4 for NeuTRENO suggested by Nguyen et al.
+(2023).
+For DenseFormer, only $\bm{\lambda}_{n,n}$ are set to 1 and the others are set to zero during initialization here.
 3.2 ResFormer
 In contrast, before Eqn.
-1, ResFormer introduces a skip connection from the first layer’s value $\mathbf{V}_{0}=\mathbf{H}_{0}\mathbf{W}^{\mathbf{V}}_{1}$ to current layer’s value $\mathbf{V}_{n}=\mathbf{H}_{n}\mathbf{W}^{\mathbf{V}}$:
-$\mathbf{V}_{n}$ $={\bm{\lambda}_{n,1}}\mathbf{V}_{1}+{\bm{\lambda}_{n,2}}\mathbf{H}_{n}\mathbf{W}^{\mathbf{V}}.$ (5)
+1, ResFormer introduces a skip connection from the first layer’s value $\mathbf{V}_{1}=\mathbf{H}_{0}\mathbf{W}^{\mathbf{V}}_{1}$ to current layer’s value $\mathbf{V}_{n}=\mathbf{H}_{n-1}\mathbf{W}^{\mathbf{V}}_{n}$:
+$\mathbf{V}_{n}$ $={\bm{\lambda}_{n,1}}\mathbf{V}_{1}+{\bm{\lambda}_{n,2}}\mathbf{H}_{n-1}\mathbf{W}^{\mathbf{V}}_{n}.$ (5)
 where ${\bm{\lambda}_{n,1}}$ and ${\bm{\lambda}_{n,2}}$ are flexible scalars.
-When all $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ are predetermined constants, it is termed Constant-ResFormer.
+When all $\bm{\lambda}_{n,1}$ and $\bm{\lambda}_{n,2}$ are predetermined constants, it is termed Constant-ResFormer.
 If $\{{\bm{\lambda}_{n,1}}={\bm{\lambda}_{n,2}}\}_{n=1}^{N}$, where $N$ is the total number of layers, the model is called Identity-ResFormer.
-Another variant where some layers have $\bm{\lambda}_{1}=0$ is referred to as Sparse-ResFormer.
-Besides, if $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ are trainable parameters, the model is termed Learnable-ResFormer.
-Unless otherwise specified, $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ are initialized to 0.5 for Learnable-ResFormer and are predetermined as 0.5 for Identity-ResFormer.
-A more general form is the Dense-ResFormer, defined as $\mathbf{V}_{n}\!=\!\bm{\lambda}_{n,n}\mathbf{H}_{n}\mathbf{W}^{\mathbf{V}}+\sum_{i=1}^{n-1}\bm{\lambda}_{n,i}\mathbf{V}_{i}$ for $n\geq 2$, where $\{\bm{\lambda}_{n}\!\}$ are constants or trainable scalars.
-Unless noted, all $\bm{\lambda}$ are initialized to 1 here.
+Another variant where some layers have $\bm{\lambda}_{n,1}=0$ is referred to as Sparse-ResFormer.
+Besides, if $\bm{\lambda}_{n,1}$ and $\bm{\lambda}_{n,2}$ are trainable parameters, the model is termed Learnable-ResFormer.
+Unless otherwise specified, $\bm{\lambda}_{n,1}$ and $\bm{\lambda}_{n,2}$ are initialized to 0.5 for Learnable-ResFormer and are predetermined as 0.5 for Identity-ResFormer.
+Furthermore, given that higher layers require more supplementary information from $\mathbf{V}_{1}$, we propose the Learnable ResFormer Plus.
+The initialization strategy is as follows: 1).
+$\bm{\lambda}_{n,2}$ is initialized to 0.5, where $n=1,2,\ldots,N$.
+2).
+$\bm{\lambda}_{n,1}$ is initialized to $\bm{\lambda_{\text{scale}}}\cdot\frac{e^{\bm{\lambda}_{n,1}^{\prime}}}{\sum_{j=1}^{N}e^{\bm{\lambda}_{j,1}^{\prime}}}$, where $\bm{\lambda_{\text{scale}}}$ is initialized to the total layer number $N$ and is shared across all layers.
+A more general form is Dense-ResFormer, defined as $\mathbf{V}_{n}\!=\!\bm{\lambda}_{n,n}\mathbf{H}_{n-1}\mathbf{W}^{\mathbf{V}}_{n}+\sum_{i=1}^{n-1}\bm{\lambda}_{n,i}\mathbf{V}_{i}$ for $n\geq 2$, where $\!\{\bm{\lambda}_{n,i}\!\}_{i=1}^{n-1}$ are constants or trainable scalars.
+Unless noted, all $\bm{\lambda}_{n,n}$ are set to 1.
 3.3 SVFormer
@@
 Table 1: Results of sharing different parts every 2 layers.
-Beyond ResFormer, SVFormer adopts standard attention in the first layer and obtain the attention output $\mathbf{U_{n}}$ for $n$-th layer where $n\geq 2$ through $\mathbf{U}_{n}=\mathbf{A}_{n}{\mathbf{V}_{1}}.$ Its main advantage is that it only requires computing and storing the value vectors for the first layer, saving nearly half of the $KV$ cache during inference.
+Figure 2: (Left) Validation loss as model size scales from 82M to 468M parameters on 20B tokens.
+(Medium) Validation loss for the 468M parameter model evaluated every 2B tokens.
+ResFormer achieves approximately 16.1%-20.3% reduction in both model parameters and training data.
+(Right) Validation loss for the 1.6B parameter model evaluated every 10B tokens.
+Model Wiki. LMB. Arc-c Arc-e BoolQ Hella. LMB. Obqa. Wino. PiQA Avg.
+PPL PPL ACC ACC ACC ACC ACC ACC ACC ACC ACC
+Transformer 24.8 33.4 18.8 49.6 57.5 31.1 34.1 17.6 49.8 66.1 40.6
+NeuTRENO 24.3 36.3 20.7 48.8 59.2 31.6 34.2 19.0 51.8 65.9 41.4
+DenseFormer 24.0 28.0 20.1 48.7 56.5 32.1 36.6 17.6 49.3 65.6 40.8
+Identity ResFormer 23.8 32.9 20.2 49.0 59.0 32.1 36.0 16.8 51.2 66.1 41.3
+Learnable ResFormer 23.7 32.5 21.2 50.3 60.9 32.3 36.3 18.8 51.2 67.0 42.3
+Learnable ResFormer plus 23.2 31.4 21.2 49.7 60.6 32.4 36.0 17.8 51.1 67.5 42.0
+Table 2: Downstream evaluation of different models with 468M parameters trained on 20B tokens.
+Beyond ResFormer, SVFormer adopts standard attention in the first layer and obtain the attention output $\mathbf{U_{n}}$ for $n$-th layer when $n\geq 2$ through $\mathbf{U}_{n}=\mathbf{A}_{n}{\mathbf{V}_{1}}$, where $\mathbf{A}_{n}$ is the attention matrix of $n$-th layer.
+Its main advantage is that it only requires computing and storing the value vectors for the first layer, saving nearly half of the $KV$ cache during inference.
 Similar methods like CLA reduce $KV$ cache by sharing both of the key and value vectors every two layers.
@@
 Specifically, the architecture includes pre-normalization, SwiGLU activations (Shazeer, 2020), rotary position embedding (Su et al., 2024), and no dropout.
-For slimpajama, we randomly sample nearly 20B tokens based to the original data distribution of seven domains during training and adopt tokenizer used for “RedPajama-INCITE-7B-Base".
+For SlimPajama, we randomly sample nearly 20B tokens based on the original data distribution of seven domains during training and adopt tokenizer used for “RedPajama-INCITE-7B-Base".
 See Table 12 in Appendix for data details.
@@
 We adopt deepspeed zero-2 optimizer and flash attention mechanism.
+Scaling Details
+We conduct scaling experiments on a 1.6B parameter model using approximately 200B tokens of internal pre-training data.
+Training is performed with a sequence length of 8,192, learning rate of 1.5e-4, and batch size of 2M tokens, requiring about 3,584 H800 GPU hours per run.
 Evaluation
@@
 Additionally, we randomly selected a fixed set of 1,000 sample sequences for subsequent visualization analysis.
-4.2 Scaling Analysis of ResFormer
-Figure 2: (Left) Validation loss as model size scales from 82M to 468M parameters.
-(Right) Validation loss for the 468M parameter model evaluated every 2B tokens.
-ResFormer achieves approximately 13.3%-15.4% reduction in both model parameters and training data.
-We analyzed how ResFormer and Transformer scale at model size and data size.
-ResFormer and Transformer are trained on similar experiment setting.
-On the one hand, we trained model with 82M, 180M, 320M and 468M parameters on 20B training tokens and evaluated them on a separate validation set.
-As shown in Fig.2 (Left), ResFormer achieves equivalent validation loss to the Transformer while utilizing 13.3% fewer model parameters.
-On the other hand, we evaluated the 468M models every 2B tokens and ResFormer needs 15.4% fewer training tokens to achieve the same loss as Transformer.
+We also compare different models on several classical reasoning tasks following (Zhang et al., 2024) in a zero-shot way.
+The tasks include Hellaswag (Hella.) (Zellers et al., 2019) Openbookqa (Obqa.) (Mihaylov et al., 2018), WinoGrande (Wino.) (Sakaguchi et al., 2019), ARC-Easy (Arc-e) and ARC-Challenge (Arc-c) (Clark et al., 2018) and PiQA (Bisk et al., 2020).
+In addition to accuracy on reasoning tasks, we also report the perplexity (PPL) on Wikitext (Wiki.) and LAMBADA (LMB.).
 4.2 ResFormer vs. NeuTRENO, DenseFormer
+We analyze how different models scale with model size and data size under similar experimental settings.
+We train models with 82M, 180M, 320M, and 468M parameters on 20B tokens and evaluate on a validation set.
+Fig.2 (Left) shows ResFormer achieves equivalent validation loss to Transformer while using 16.11% fewer parameters.
+We also evaluate 468M models every 2B tokens, finding ResFormer requires 20.3% fewer training tokens to match Transformer’s loss (Fig.2 (Medium)).
+All ResFormer variants demonstrate superior scaling compared to NeuTRENO and DenseFormer.
+Table 2 confirms ResFormer outperforms other models on downstream tasks, with Learnable ResFormer achieving 1.7 point average accuracy improvement over vanilla Transformer.
+To validate the performance improvement of value residual at larger training scales, we conduct additional experiments following the scaling details in Sec.4.1.
+As shown in Fig.2 (Right), value residual consistently improves performance throughout training.
+Figure 3: The impact of varying $\bm{\lambda}$ values on 82M 8-layer Constant-ResFormer and NeuTRENO.
+Both Constant-ResFormer and NeuTRENO rely on predetermined $\bm{\lambda}$ constants.
+Fig.
+3 shows the performance curves of these models against varying $\bm{\lambda}$.
+Results indicate that Constant-ResFormer significantly outperforms NeuTRENO and demonstrates greater robustness across a wider range of $\bm{\lambda}$ values, achieving optimal performance at $\bm{\lambda}=2$.
 Model Initial Form Loss
@@
 Vanilla Transformer - 2.739
-$\text{DenseFormer}^{*}$ $\mathbf{H}_{n}^{\prime}\!=\!{\color[rgb]{1,0,0}1\times}\mathbf{H}_{i}+\sum_{i=1}^{n-1}{\color[rgb]{1,0,0}0\times}\mathbf{H}_{i}$ 2.722
+$\text{DenseFormer}^{*}$ $\mathbf{H}_{n}^{\prime}\!=\!{\color[rgb]{1,0,0}1\times}\mathbf{H}_{n}+\sum_{i=0}^{n-1}{\color[rgb]{1,0,0}0\times}\mathbf{H}_{i}$ 2.722
 NeuTRENO $\mathbf{U}_{n}^{\prime}={\color[rgb]{1,0,0}0.4}(\mathbf{V}_{1}-\mathbf{V}_{n})+\mathbf{U}_{n}$ 2.72
@@
 […]
-Constant-ResFormer $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}2}\mathbf{V}_{1}+{\color[rgb]{1,0,0}0.5}\mathbf{V}_{n}$ 2.7
+Constant-ResFormer $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}2}\mathbf{V}_{1}+\mathbf{V}_{n}$ 2.7
 Sparse-ResFormer $\left\{\begin{array}[]{l}\mathbf{V}_{n}^{\prime}=\mathbf{V}_{n},1\leq n\leq 5\\ \mathbf{V}_{n}^{\prime}=\mathbf{V}_{1},6\leq n\leq 8\end{array}\right.$ 2.696
-Sparse-ResFormer $\left\{\begin{array}[]{l}\mathbf{V}_{n}^{\prime}=\mathbf{V}_{n},1\leq n\leq 5\\ \mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}5}\mathbf{V}_{1}+{\color[rgb]{1,0,0}0.5}\mathbf{V}_{n},\\ \qquad\ \quad\qquad 6\leq n\leq 8\end{array}\right.$ 2.687
+Sparse-ResFormer $\left\{\begin{array}[]{l}\mathbf{V}_{n}^{\prime}=\mathbf{V}_{n},1\leq n\leq 5\\ \mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}5}\mathbf{V}_{1}+\mathbf{V}_{n},\\ \qquad\ \quad\qquad 6\leq n\leq 8\end{array}\right.$ 2.687
+[…]
 Table 3: Average valid loss for 8-layer, 82M-parameter models.
@@
 For models marked with “*" , $\bm{\lambda}$ is learnable, and the red numbers indicate the initial value; otherwise, red numbers are fixed constants.
-We compared the average validation loss of different models with 82M parameters trained on 20B tokens.
-In our experimental setup, NeuTRENO performed best with $\bm{\lambda}=0.4$, achieving comparable results to DenseFormer.
-All ResFormer variants, including the simplest Identity-ResFormer, demonstrated significant performance improvements.
+Furthermore, we test the performance of other varaints of ResFormer mentioned in Sec.3.2.
 The $\bm{\lambda}$ values for Constant-ResFormer and Sparse-ResFormer were optimized through multiple experiments.
-While Sparse-ResFormer achieved the best overall performance, it is challenging to pre-determine the optimal layers for $\mathbf{V}_{1}$ connections and their corresponding $\bm{\lambda}_{1}$ values in more general scenarios.
-Although, Learnable-ResFormers are struggle to identify the optimal value residual pattern, they outperform Identity-ResFormer.
-Interestingly, the second to last row shows that Sparse-ResFormer achieved better performance despite having three fewer $\mathbf{W}^{\mathbf{V}}$.
-Figure 3: The impact of varying $\bm{\lambda}$ values on 82M 8-layer Constant-ResFormer and NeuTRENO.
-Both Constant-ResFormer and NeuTRENO rely on predetermined $\bm{\lambda}$ constants.
-Fig.
-3 shows the performance curves of these models against varying $\bm{\lambda}$.
-Results indicate that Constant-ResFormer significantly outperforms NeuTRENO and demonstrates greater robustness across a wider range of $\bm{\lambda}$ values, achieving optimal performance at $\bm{\lambda}=2$.
+All ResFormer variants, including the simplest Identity-ResFormer, show significant performance improvements.
+However, manually-tuned Sparse-ResFormer and Learnable-ResFormer-Plus outperform the standard Learnable-ResFormer.
+It demonstrates the challenge to pre-determine the optimal layers for $\mathbf{V}_{1}$ connections and their corresponding $\bm{\lambda}_{1}$ values in more general scenarios.
+Interestingly, the third to last row shows that Sparse-ResFormer achieved better performance despite having three fewer $\mathbf{W}^{\mathbf{V}}$.
 4.3 Truly Better or Just Faster?
@@
 - - 2.7389
-value residual $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}0.5}\mathbf{V}_{1}+{\color[rgb]{1,0,0}0.5}\mathbf{V}_{n}$ 2.7049
-hidden residual $\mathbf{H}_{n}^{\prime}={\color[rgb]{1,0,0}0.5}\mathbf{H}_{0}+{\color[rgb]{1,0,0}0.5}\mathbf{H}_{n}$ 2.7812
-value residual $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}0\times}\mathbf{V}_{1}+{\color[rgb]{1,0,0}1}\mathbf{V}_{n}$ 2.7298
-hidden residual $\mathbf{H}_{n}^{\prime}={\color[rgb]{1,0,0}0\times}\mathbf{H}_{0}+{\color[rgb]{1,0,0}1}\mathbf{H}_{n}$ 2.7216
+value $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}0.5}\mathbf{V}_{1}+{\color[rgb]{1,0,0}0.5}\mathbf{V}_{n}$ 2.705
+hidden $\mathbf{H}_{n}^{\prime}={\color[rgb]{1,0,0}0.5}\mathbf{H}_{0}+{\color[rgb]{1,0,0}0.5}\mathbf{H}_{n}$ 2.781
+value $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}0\times}\mathbf{V}_{1}+{\color[rgb]{1,0,0}1}\mathbf{V}_{n}$ 2.73
+hidden $\mathbf{H}_{n}^{\prime}={\color[rgb]{1,0,0}0\times}\mathbf{H}_{0}+{\color[rgb]{1,0,0}1}\mathbf{H}_{n}$ 2.722
 Table 4: Comparison of additional value residual (to $\mathbf{V}_{1}$) and hidden residual (to $\mathbf{H}_{0}$) connections against the default hidden residual, under various $\mathbf{\lambda}$ initializations.
 Trainable $\mathbf{\lambda}$ parameters are highlighted in red.
-Our experiments revealed that $\mathbf{V}_{1}$ information provides additional benefits to later network layers, despite both $\mathbf{H}_{1}$ and $\mathbf{V}_{1}$ containing initial, unfused token information.
+Our experiments revealed that $\mathbf{V}_{1}$ information provides additional benefits to later network layers, despite both $\mathbf{H}_{0}$ and $\mathbf{V}_{1}$ containing initial, unfused token information.
 $\mathbf{H}_{0}$ is propagated through default hidden residual connections, but it may be diluted by subsequent information, hindering its effective utilization in later layers.
-To test this hypothesis, we introduced an additional skip connection to $\mathbf{H}_{0}$ : $\mathbf{H}_{n}^{\prime}={\mathbf{\lambda}_{1}}\mathbf{H}_{1}+{\mathbf{\lambda}_{2}}\mathbf{H}_{n}$, where $\mathbf{\lambda}$ is learnable.
+To test this hypothesis, we introduced an additional skip connection to $\mathbf{H}_{0}$ : $\mathbf{H}_{n}^{\prime}={\mathbf{\lambda}_{0}}\mathbf{H}_{1}+{\mathbf{\lambda}_{2}}\mathbf{H}_{n}$, where $\mathbf{\lambda}$ is learnable.
 We conducted experiments with two $\mathbf{\lambda}$ initialization settings and compared them to value residual.
@@
 Hidden Residual Starts Place Value Residual Target Place Valid Loss
-$\mathbf{H_{0}}$ (Default) - 2.7389
-$\mathbf{H_{0}}$ (Default) $\mathbf{V_{1}}$ 2.7117
-$\mathbf{H_{0}}$ (Default) $\mathbf{V_{2}}$ 2.7375
-$\mathbf{H_{1}}$ 2.7802
-$\mathbf{H_{1}}$ $\mathbf{V_{2}}$ 2.784
-$\mathbf{H_{2}}$ 2.8196
-$\mathbf{H_{2}}$ $\mathbf{V_{2}}$ 2.7873
-$\mathbf{H_{2}}$ 2.8196
-$\mathbf{H_{2}}$ $\mathbf{V_{3}}$ 2.8333
-$\mathbf{H_{3}}$ 3.0568
-$\mathbf{H_{3}}$ $\mathbf{V_{3}}$ 2.8832
+$\mathbf{H_{0}}$ (Default) - 2.739
+$\mathbf{H_{0}}$ (Default) $\mathbf{V_{1}}$ 2.712
+$\mathbf{H_{0}}$ (Default) $\mathbf{V_{2}}$ 2.738
+$\mathbf{H_{1}}$ 2.78
+$\mathbf{H_{1}}$ $\mathbf{V_{2}}$ 2.78
+$\mathbf{H_{2}}$ 2.82
+$\mathbf{H_{2}}$ $\mathbf{V_{2}}$ 2.787
+$\mathbf{H_{2}}$ 2.82
+$\mathbf{H_{2}}$ $\mathbf{V_{3}}$ 2.833
+$\mathbf{H_{3}}$ 3.057
+$\mathbf{H_{3}}$ $\mathbf{V_{3}}$ 2.883
 Table 5: Comparison of performance across different value residual target and varying hidden residual start settings.
@@
 4.6 SVFormer vs. GQA,CLA
-Sequence Length Model Valid Loss
+Sequence Model Valid
+Length Loss
 2,048 - 2.739
@@
 SVFormer 2.7485
-Sequence Length Model Valid Loss
+Sequence Model Valid
+Length Loss
 64,000 - 2.753
@@
 Right: Analysis of critical point, and we predict it for length 64,000 using linear regression with the last 1,000 data points.
-Architecture Learning Rate Warnup Steps Valid Loss
-Llama (82M) 1e-4 120 -0.033
+Model Type Learning Rate Warnup Steps Valid Loss
+Llama 1e-4 120 -0.033
 3e-4 120 +0.021
@@
 6e-4 1,200 +0.036
-GPT2 (78M) 6e-4 120 +0.029
+GPT2 6e-4 120 +0.029
 Table 9: Relative validation loss of SVFormer compared to vanilla Transformer under different hyper-parameter settings.
@@
 On the other hand, the models trained in our study are solely utilized for experimental analysis and will not be publicly deployed.
+Acknowledgments
+This work was supported by the Scientific Research Project of Westlake University (Grant No. WU2024B003).
 Appendix A Appendix
@@
 This further suggests that the value residual patterns learned by the Learnable ResFormer do not introduce significant negative effects in this context.
-A.2 Downstream Evaluations
-We compare the ResFormer and vanilla Transformer on several classical reasoning tasks following (Zhang et al., 2024) in a zero-shot way.
-The tasks include Hellaswag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2019), ARC-Easy and ARC-Challenge (Clark et al., 2018) and PIQA (Bisk et al., 2020).
-The results in Table 11 show that ResFormer achieved an average accuracy improvement of nearly 3% compared to the vanilla Transformer.
-Model Max Length HellaSwag Obqa WinoGrande ARC-c ARC-e PIQA Avg
-Transformer 2,048 0.263 0.142 0.492 0.199 0.331 0.572 0.333
-ResFormer 2,048 0.273 0.148 0.512 0.182 0.414 0.604 0.355
-Transformer 64,000 0.267 0.142 0.485 0.179 0.322 0.570 0.328
-ResFormer 64,000 0.274 0.136 0.513 0.184 0.407 0.588 0.350
-Table 11: Zero-shot accuracy on commonsense reasoning tasks.
-A.3 Pre-train Dataset
+A.2 Pre-train Dataset
 Data source proportions Tokens
@@
 In this way, we sampled 20B tokens data from original 627B data based on the original data proportions shown in the Table 12.
-A.4 Training Details
+A.3 Training Details
 Max Sequence Length 512 2,048 8,192 32,000 64,000
@@
 (2023).
-Additionally, the $\bm{\lambda}$ in Eqn.
-3 is set to be 0.4 for NeuTRENO.
 Moreover, as reported in Table 13, the batch size that a single GPU can accommodate varies depending on the length of the training sequences.